### v1

# On local Galois deformation rings: generalised tori

Abstract.
We study deformation theory of mod $p$ Galois representations of $p$-adic fields with values in generalised tori, such as $L$-groups of (possibly non-split) tori. We show that the corresponding deformation rings are formally smooth over a group algebra of a finite abelian $p$-group. We compute their dimension and the set of irreducible components.
1. Introduction
Let $p$ denote any prime number, let $F$ be a finite extension of $\mathbb{Q}_{p}$, and let $\Gamma_{F}$ denote its absolute Galois group. Let $L$ be another finite extension of $\mathbb{Q}_{p}$ with ring of integers $\mathcal{O}$, uniformiser $\varpi$ and residue field $k=\mathcal{O}/\varpi$.
Let $G$ be a smooth affine group scheme over $\mathcal{O}$, such that its neutral component $G^{0}$ is torus and the component group $G/G^{0}$ is a finite étale group scheme over $\mathcal{O}$. We call such group schemes generalised tori. We do not make any assumptions regarding $p$ and the component group of $G$.
After replacing $L$ by a finite unramified extension one may assume that $G^{0}$ is a split torus and $G/G^{0}$ is a constant group scheme. We will assume this for the rest of the introduction.
An important example of a generalised torus that we have in mind is the $L$-group of a torus defined over $F$: let $H$ be a torus over $F$, which splits over a finite Galois extension $E$ over $F$, then the $L$-group of $H$ is ${}^{L}H=\hat{H}\rtimes\underline{\operatorname{Gal}(E/F)}$, where $\hat{H}$ is the split $\mathcal{O}$-torus, such that the character lattice of $\hat{H}$ is equal to the cocharacter lattice of $H_{E}$ and $\underline{\operatorname{Gal}(E/F)}$ is the constant group scheme associated to $\operatorname{Gal}(E/F)$. In this example, the surjection $G\twoheadrightarrow G/G^{0}$ has a section of group schemes. We do not assume this in general.
We fix a continuous representation $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ and denote by $D^{\square}_{\overline{\rho}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ the functor from the category $\mathfrak{A}_{\mathcal{O}}$ of local artinian $\mathcal{O}$-algebras with residue field $k$ to the category of sets, such that for $(A,\mathfrak{m}_{A})\in\mathfrak{A}_{\mathcal{O}}$, $D^{\square}_{\overline{\rho}}(A)$ is the set of continuous representations $\rho_{A}:\Gamma_{F}\rightarrow G(A)$, such that
$\rho_{A}(\gamma)\equiv\overline{\rho}(\gamma)\pmod{\mathfrak{m}_{A}},\quad\forall\gamma\in\Gamma_{F}.$
The functor $D^{\square}_{\overline{\rho}}$ of framed deformations of $\overline{\rho}$ is pro-represented by a complete local noetherian $\mathcal{O}$-algebra $R^{\square}_{\overline{\rho}}$ with residue field $k$.
Theorem 1.1.
There is a finite extension $L^{\prime}$ of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and a continuous representation $\rho:\Gamma_{F}\rightarrow G(\mathcal{O}^{\prime})$ lifting $\overline{\rho}$.
Remark 1.2.
If $G\cong G^{0}\rtimes(G/G^{0})$ then a lift in Theorem 1.1 can be constructed over $\mathcal{O}$ using the Teichmüller lift to define a section of $G^{0}(\mathcal{O})\twoheadrightarrow G^{0}(k)$, which then induces a section $\sigma:G(k)\rightarrow G(\mathcal{O})$. Then $\sigma\circ\overline{\rho}$ is the required lift. In particular, in this case Theorem 1.1 holds if we replace $\Gamma_{F}$ with any profinite group. However, it seems non-trivial to prove 1.1 in general and our argument uses that the Euler–Poincaré characteristic formula holds for $\Gamma_{F}$.
Let $\Gamma_{E}$ be the kernel of the composition $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\rightarrow(G/G^{0})(k)$ and let $\Delta$ be the image of this map. We identify $\Delta=\operatorname{Gal}(E/F)$. Let $M$ be the character lattice of $G^{0}$. The action of $G$ on $G^{0}$ by conjugation induces an action of $\Delta$ on $M$. Let $\Gamma_{E}^{\mathrm{ab},p}$ be the maximal abelian pro-$p$ quotient of $\Gamma_{E}$. Below we will consider the diagonal action of $\Delta$ on $\Gamma_{E}^{\mathrm{ab},p}\otimes M$ and $\mu_{p^{\infty}}(E)\otimes M$, where $\mu_{p^{\infty}}(E)$ is the subgroup of $p$-power roots of unity in $E$.
Theorem 1.3.
Assume that Theorem 1.1 holds with $\mathcal{O}=\mathcal{O}^{\prime}$ then
$R^{\square}_{\overline{\rho}}\cong\mathcal{O}[(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}]\llbracket x_{1},\ldots,x_{m}\rrbracket,$
where $m=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot([F:\mathbb{Q}_{p}]+1)$.
Let $\overline{\Theta}$ be the $G$-pseudocharacter associated to $\overline{\rho}$ and let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the deformation functor such that $D^{\mathrm{ps}}_{\overline{\Theta}}(A)$ is the set of continuous $A$-valued $G$-pseudocharacters deforming $\overline{\Theta}$. These notions are reviewed in Section 6. The functor $D^{\mathrm{ps}}_{\overline{\Theta}}$ is pro-represented by a complete local noetherian $\mathcal{O}$-algebra $R^{\mathrm{ps}}_{\overline{\Theta}}$ with residue field $k$ and we denote the universal deformation by $\Theta^{u}$. Sending a deformation of $\overline{\rho}$ to its $G$-pseudocharacter induces a natural transformation $D^{\square}_{\overline{\rho}}\rightarrow D^{\mathrm{ps}}_{\overline{\Theta}}$, and hence a map of local $\mathcal{O}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$. If $G$ is a torus or more generally, when $G$ is commutative, then representations and $G$-pseudocharacters coincide and this map is an isomorphism.
Theorem 1.4.
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ is formally smooth. Moreover, if Theorem 1.1 holds with $\mathcal{O}^{\prime}=\mathcal{O}$ then
$R^{\mathrm{ps}}_{\overline{\Theta}}\cong\mathcal{O}\llbracket(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}\rrbracket\cong\mathcal{O}[(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}]\llbracket x_{1},\ldots,x_{r}\rrbracket,$
where $r=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot[F:\mathbb{Q}_{p}]+\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
We let $X^{\mathrm{gen}}_{G,\overline{\Theta}}:R^{\mathrm{ps}}_{\overline{\Theta}}\text{-}\mathrm{alg}\rightarrow\operatorname{Set}$ be the functor such that $X^{\mathrm{gen}}_{G,\overline{\Theta}}(A)$ is the set of representations $\rho:\Gamma_{F}\rightarrow G(A)$ such that its $G$-pseudocharacter $\Theta_{\rho}$ satisfies $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A$. This functor is representable by a finite type $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra $A^{\mathrm{gen}}_{G,\overline{\Theta}}$.
Theorem 1.5.
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is smooth. Moreover, if Theorem 1.1 holds with $\mathcal{O}^{\prime}=\mathcal{O}$ then
$A^{\mathrm{gen}}_{G,\overline{\Theta}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
where $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
We note that $\mu_{p^{\infty}}(E)$ is a finite $p$-group. This implies that $\mu:=(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$ is also a finite $p$-group and we denote its order by $p^{m}$. We assume further that $\mathcal{O}$ contains all $p^{m}$-th roots of $1$ and let $\mathrm{X}(\mu)$ be the group of characters $\chi:\mu\rightarrow\mathcal{O}^{\times}$. It then follows from Theorems 1.3, 1.4, 1.5 that the sets of irreducible components of $R^{\square}_{\overline{\rho}}$, $R^{\mathrm{ps}}_{\overline{\Theta}}$ and $A^{\mathrm{gen}}_{G,\overline{\Theta}}$, respectively, are in $\mathrm{X}(\mu)$-equivariant bijection with $\mathrm{X}(\mu)$. Moreover, the irreducible components and their special fibres are regular. The identification of the set of components with $\mathrm{X}(\mu)$ is non-canonical in general, as one has to distinguish one component, which corresponds to the trivial character. However, we explain in Section 7.4 that there is a canonical action of $\mathrm{X}(\mu)$ on the set of irreducible components, which is faithful and transitive.
Remark 1.6.
Let us point out that if $G\cong G^{0}\rtimes G/G^{0}$ then the lift constructed in 1.2 is canonical, as it is minimally ramified. This distinguishes the irreducible component that it lies on. Thus in this case there is a canonical $\mathrm{X}(\mu)$-equivariant bijection between $\mathrm{X}(\mu)$ and the set of irreducible components.
The above theorems are used in an essential way in a companion paper [11], where we study deformations of Galois representations with values in generalised reductive $\mathcal{O}$-group schemes $G$, which means that the neutral component $G^{0}$ is reductive and the component group $G/G^{0}$ is finite étale. If we let $G^{\prime}$ be the derived subgroup scheme of $G^{0}$, then $G^{0}/G^{\prime}$ is a torus and $G/G^{\prime}$ is a generalised torus as considered in this paper. If $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ is a continuous representation then we show in [11, Proposition 4.20] that its deformation ring $R^{\square}_{\overline{\rho}}$ can be presented over the deformation ring of $\varphi\circ\overline{\rho}:\Gamma_{F}\rightarrow(G/G^{\prime})(k)$, where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map. This allows us to split up the arguments into “torus part”, carried out in this paper, and “semisimple part” carried out in [11].
If $\mu$ is any finite abelian $p$-group then $\mu\cong\prod_{i=1}^{l}\mathbb{Z}/p^{e_{i}}$ and hence
$\mathcal{O}[\mu]\cong\bigotimes_{i=1}^{l}\frac{\mathcal{O}[z_{i}]}{((1+z_{i})^{p^{e_{i}}}-1)}$
is complete intersection. It then follows from Theorems 1.3, 1.4, 1.5 that the rings $R^{\square}_{\overline{\rho}}$, $R^{\mathrm{ps}}_{\overline{\Theta}}$ and $A^{\mathrm{gen}}_{G,\overline{\Theta}}$ are locally complete intersections. To relate their dimensions to the dimensions of the deformation rings appearing in [11] we note that $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M=\dim G_{k}$ and, if $\Delta=(G/G^{0})(\bar{k})$ (which we may assume without changing the functors $D^{\square}_{\overline{\rho}}$, $D^{\mathrm{ps}}_{\overline{\Theta}}$ and $X^{\mathrm{gen}}_{G,\overline{\Theta}}$), then $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}=\dim Z(G)_{k}$. The scheme $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ coincides with the scheme denoted by $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ in [11], when $G$ is a generalised torus. In [11] in the definition of $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ an additional condition is imposed on representations, called $R^{\mathrm{ps}}_{\overline{\Theta}}$-condensed, and we show in Lemma 9.6, that when $G$ is a generalised torus, this condition holds automatically.
1.1. Arithmetic setting
Let us first recall how one proves the result when $G=\mathbb{G}_{m}$. The first step is to observe that Theorem 1.1 holds with $\mathcal{O}^{\prime}=\mathcal{O}$ for example using the Teichmüller lift $\sigma:k^{\times}\rightarrow\mathcal{O}^{\times}$ and letting $\rho_{0}=\sigma\circ\overline{\rho}$. The second step is to observe that the map $\rho\mapsto\rho\rho^{-1}_{0}$ induces a bijection between $D^{\square}_{\overline{\rho}}\rightarrow D^{\square}_{\mathbf{1}}$, where $\mathbf{1}$ is the trivial representation. One may identify $D^{\square}_{\mathbf{1}}(A)$ with the set of continuous group homomorphisms $\Gamma_{F}\rightarrow 1+\mathfrak{m}_{A}$ and since the target is an abelian $p$-group the functor is representable by $\mathcal{O}\llbracket\Gamma_{F}^{\mathrm{ab},p}\rrbracket$. An interesting feature of this argument is that $D^{\square}_{\mathbf{1}}$ takes values in $\operatorname{Ab}$ as opposed to $\operatorname{Set}$ and the map
$D_{\mathbf{1}}^{\square}(A)\times D_{\overline{\rho}}^{\square}(A)\rightarrow D^{\square}_{\overline{\rho}}(A),\quad(\Phi,\rho)\mapsto\Phi\rho$
defines an action of an abelian group $D_{\mathbf{1}}^{\square}(A)$ on the set $D^{\square}_{\overline{\rho}}(A)$. Moreover, if the set is non-empty then this action is faithful and transitive.
If $G$ is connected then $G=\mathfrak{D}(M)$, where $M$ is the character lattice, and so $G(A)=\operatorname{Hom}(M,A^{\times})$. The same argument goes through and we only need to observe that
$\operatorname{Hom}^{\mathrm{cont}}_{\mathrm{Group}}(\Gamma_{F},G(A))\cong\operatorname{Hom}^{\mathrm{cont}}_{\mathrm{Group}}(\Gamma_{F}^{\mathrm{ab}}\otimes M,A^{\times}),$
which implies that $D^{\square}_{\overline{\rho}}$ is representable by $\mathcal{O}\llbracket\Gamma_{F}^{\mathrm{ab},p}\otimes M\rrbracket$.
Let us consider the general case now. The first issue, alluded to in 1.2, is that it is not clear anymore that such a lift exists. However, let us postpone the explanation how we deal with this problem until Section 1.3 and let us give ourselves a lift $\rho_{0}:\Gamma_{F}\rightarrow G(\mathcal{O})$ of $\overline{\rho}$. Then the map $\rho\mapsto[\gamma\mapsto\rho(\gamma)\rho_{0}(\gamma)^{-1}]$ considered above induces a bijection between $D^{\square}_{\overline{\rho}}(A)$ and the set $\widehat{Z}^{1}(A)$ of continuous $1$-cocycles $\Phi:\Gamma_{F}\rightarrow\operatorname{Hom}(M,1+\mathfrak{m}_{A})$. Further, one may interpret $\widehat{Z}^{1}$ as $D^{\square}_{\overline{\rho}}$ for the group $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ given by $\overline{\rho}(\gamma)=(1,\pi(\gamma))$, where $\pi:\Gamma_{F}\rightarrow\Delta$ is the quotient map. So that $\widehat{Z}^{1}$ is the analogue of $D^{\square}_{\mathbf{1}}$ and we have to find the ring pro-representing the functor $\widehat{Z}^{1}$.
Before we describe our solution, let us point out one dead end. If $\Phi\in\widehat{Z}^{1}(A)$ then its restriction to $\Gamma_{E}$ is just a continuous homomorphism
$\Phi|_{\Gamma_{E}}:\Gamma_{E}\rightarrow\operatorname{Hom}(M,1+\mathfrak{m}_{A}),$
as $\Gamma_{E}$ acts trivially on $M$. This induces a continuous homomorphism
$\varphi:\Gamma_{E}^{\mathrm{ab},p}\otimes M\rightarrow 1+\mathfrak{m}_{A},$
and the fact that the homomorphism is obtained by a restriction of a cocycle to $\Gamma_{E}$ implies that $\varphi$ factors through $(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}\rightarrow 1+\mathfrak{m}_{A}$. However, in Theorem 1.3 we have $\Delta$-invariants and not $\Delta$-coinvariants appearing. Tate group cohomology gives us an exact sequence
$0\rightarrow\widehat{H}^{-1}(\Delta,\Gamma_{E}^{\mathrm{ab},p}\otimes M)\rightarrow(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}\rightarrow(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}\rightarrow\widehat{H}^{0}(\Delta,\Gamma_{E}^{\mathrm{ab},p}\otimes M)\rightarrow 0$
One may show that the Tate group cohomology groups $\widehat{H}^{i}(\Delta,\Gamma_{E}^{\mathrm{ab},p}\otimes M)$ are finite for all $i\in\mathbb{Z}$, but there is no reason why they should vanish, unless $p$ does not divide $|\Delta|$. We conclude that the restriction to $\Gamma_{E}$ will lose information in general.
Our solution is motivated by Langlands paper [10] on his correspondence for tori, where he shows that the universal coefficient theorem induces an isomorphism
(1) $H^{1}(W_{E/F},\operatorname{Hom}(M,\mathbb{C}^{\times}))\cong\operatorname{Hom}(H_{1}(W_{E/F},M),\mathbb{C}^{\times}))$
and there are natural isomorphisms
(2) $H_{1}(W_{E/F},M)\cong H_{1}(E^{\times},M)^{\Delta}\cong(E^{\times}\otimes M)^{\Delta},$
where the Weil group $W_{E/F}$ fits into a short exact sequence
(3) $0\rightarrow E^{\times}\rightarrow W_{E/F}\rightarrow\Delta\rightarrow 0,$
corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$. Since every continuous representation $\Gamma_{F}\rightarrow G(A)$ factors through the profinite completion of $W_{E/F}$, this motivates us to study the functor
$\operatorname{CRing}\rightarrow\operatorname{Ab},\quad A\mapsto Z^{1}(W_{E/F},G(A))$
without imposing any continuity conditions on the cocycles and allowing any commutative ring $A$. We determine the ring representing this functor and show that the ring representing $\widehat{Z}^{1}$ arises as a completion of it with respect to the maximal ideal corresponding to the trivial cocycle.
1.2. Abstract setting
We carry out this study in Section 4 in the following abstract setting: let $\Gamma_{1}$ be an abstract group, $\Gamma_{2}$ a normal subgroup of $\Gamma_{1}$ of finite index and let $\Delta:=\Gamma_{1}/\Gamma_{2}$. The exact sequence
(4) $0\rightarrow\Gamma_{2}^{\mathrm{ab}}\rightarrow\Gamma_{1}/[\Gamma_{2},\Gamma_{2}]\rightarrow\Delta\rightarrow 0$
defines a class in $H^{2}(\Delta,\Gamma_{2}^{\mathrm{ab}})$. Its image in $\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})$, where $I_{\Delta}$ is the augmentation ideal in the group ring $\mathbb{Z}[\Delta]$ defines an extension of $\mathbb{Z}[\Delta]$-modules
(5) $0\rightarrow\Gamma_{2}^{\mathrm{ab}}\rightarrow\mathcal{E}\rightarrow I_{\Delta}\rightarrow 0$
Theorem 1.7.
The functor
$\operatorname{CRing}\rightarrow\operatorname{Ab},\quad A\mapsto Z^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times}))$
is represented by the group algebra $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]$.
The fpqc sheafification of the functor
$\operatorname{CRing}\rightarrow\operatorname{Ab},\quad A\mapsto H^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times}))$
is representable by the group algebra $\mathbb{Z}[H_{1}(\Gamma_{1},M)]$.
We prove 1.7 by first replacing $M$ with $\mathbb{Z}[\Delta]\otimes M$. In this case
$\operatorname{Hom}(\mathbb{Z}[\Delta]\otimes M,A^{\times})\cong\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,A^{\times})$
and we can use Shapiro’s lemma to identify $H^{1}$ with
$\operatorname{Hom}(\Gamma_{2},\operatorname{Hom}(M,A^{\times}))\cong\operatorname{Hom}(\Gamma_{2}^{\mathrm{ab}}\otimes M,A^{\times})$
Further, after choosing a representative $\overline{c}$ for each coset $c\in\Gamma_{1}/\Gamma_{2}$, we construct in Section 3 a homomorphism of abelian groups,
$H^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}V)\rightarrow Z^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}V),$
for any $\Gamma_{2}$-module $V$, which is functorial in $V$ and becomes the identity, when composed with the natural quotient map. These ingredients allow us to show that the functor
(6) $\operatorname{Ab}\rightarrow\operatorname{Ab},\quad V\mapsto Z^{1}(\Gamma_{1},\operatorname{Hom}(\mathbb{Z}[\Delta]\otimes M,V))$
is representable by $(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta})\otimes M$. We then define an action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$, which commutes with the action of $\Gamma_{1}$, such that
$Z^{1}(\Gamma_{1},\operatorname{Hom}(M,V))\cong Z^{1}(\Gamma_{1},\operatorname{Hom}(\mathbb{Z}[\Delta]\otimes M,V))^{\Delta}.$
To prove 1.7 we verify that the action of $\Delta$ on $(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta})\otimes M$ induced by its action on the functor in (6) is isomorphic to $\mathcal{E}\otimes M$. This calculation is carried out in Sections 4. The last part of Theorem 1.7 follows by observing that (1) remains an isomorphism if $\mathbb{C}^{\times}$ is replaced by any divisible group and for any $A$ there is a faithfully flat map $A\rightarrow B$ such that $B^{\times}$ is divisible.
Let $G$ be a generalised torus over some base ring $\mathcal{O}$. Let $\mathrm{PC}^{\Gamma_{1}}_{G}:\mathcal{O}\text{-}\mathrm{alg}\rightarrow\operatorname{Set}$ be the functor such that $\mathrm{PC}^{\Gamma_{1}}_{G}(A)$ is the set of $A$-valued $G$-pseudocharacters of $\Gamma_{1}$. Let $\operatorname{Rep}^{\Gamma_{1}}_{G}:\mathcal{O}\text{-}\mathrm{alg}\rightarrow\operatorname{Set}$ be the functor such that $\operatorname{Rep}^{\Gamma_{1}}_{G}(A)$ is the set of all representations $\rho:\Gamma_{1}\rightarrow G(A)$.
Theorem 1.8.
Mapping a representation to its $G$-pseudocharacter induces an isomorphism of schemes
$\operatorname{Rep}^{\Gamma_{1}}_{G}\sslash G^{0}\overset{\cong}{\longrightarrow}\mathrm{PC}^{\Gamma_{1}}_{G}$
The proof of 1.8 follows the arguments of Emerson–Morel [7], where they prove an analogous result, when $G$ is a connected reductive group over a field of characteristic 00. Since in our case $G^{0}$ is a torus, it is linearly reductive over $\mathcal{O}$, and their arguments carry over integrally.
Let us assume that $\operatorname{Rep}^{\Gamma_{1}}_{G}(\mathcal{O})$ is non-empty, $G^{0}$ is split over $\mathcal{O}$ and $G/G^{0}$ is a constant group scheme. Let $\rho_{0}:\Gamma_{1}\rightarrow G(\mathcal{O})$ be a representation and let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ be the subfunctor of $\operatorname{Rep}^{\Gamma_{1}}_{G}$ consisting of representations $\rho$ such that $\Pi\circ\rho=\Pi\circ\rho_{0}$, where $\Pi:G\rightarrow G/G^{0}$. One may similarly define $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$. We show that the map
$Z^{1}(\Gamma_{1},G^{0}(A))\rightarrow\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A),\quad\Phi\mapsto\Phi\rho_{0}$
is bijective and use Theorem 1.7 to prove:
Theorem 1.9.
The functors $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ and $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$ are represented by the group algebras $\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ and $\mathcal{O}[H_{1}(\Gamma_{1},M)]$, respectively.
If $\Gamma_{1}=W_{E/F}$ then $H_{1}(\Gamma_{1},M)\cong(E^{\times}\otimes M)^{\Delta}$ by (2). To prove 1.4 we relate the completions of the local rings of $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ and $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$ to deformation rings parameterising continuous deformations and $G$-pseudocharacters of the profinite completion $\widehat{\Gamma}_{1}$ of $\Gamma_{1}$. These arguments are carried out in Section 7.
We expect that the abstract setting will also be useful studying deformations of global Galois groups.
1.3. Producing a lift
Let us go back to the problem of exhibiting a lift of $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ to the characteristic zero as claimed in Theorem 1.1. Using $\overline{\rho}$ instead of $\rho_{0}$, we may identify the restriction of $D^{\square}_{\overline{\rho}}$ to $\mathfrak{A}_{k}$ with the restriction of $\widehat{Z}^{1}$ to $\mathfrak{A}_{k}$. This allows us to conclude that
$R^{\square}_{\overline{\rho}}/\varpi\cong k\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}\rrbracket,$
where $\mathcal{E}$ is defined by (5), where (4) is equal to (3), so that $\Gamma_{1}=W_{E/F}$ and $\Gamma_{2}=E^{\times}$ and $(-)^{\wedge,p}$ denotes the pro-$p$ completion.
In Section 8 we compute
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot([F:\mathbb{Q}_{p}]+1),$
which is then also equal to $\dim R^{\square}_{\overline{\rho}}/\varpi$. On the other hand, using Mazur’s obstruction theory and the Euler–Poincaré characteristic formula we show that
$R^{\square}_{\overline{\rho}}\cong\frac{\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket}{(f_{1},\ldots,f_{s})},$
where $r-s=\dim_{k}\operatorname{Lie}G_{k}\cdot([F:\mathbb{Q}_{p}]+1)$. Since $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M=\dim_{k}\operatorname{Lie}G_{k}$, a little commutative algebra implies that $\varpi,f_{1},\ldots,f_{s}$ is a regular sequence in $\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket$ and hence $R^{\square}_{\overline{\rho}}$ is $\mathcal{O}$-torsion free. A closed point of the generic fibre gives us the required lift.
1.4. Overview by section
In Section 2 we define what a generalised torus is. In Section 3 we establish an explicit version of Shapiro’s lemma. In Section 4 we compute the ring representing the functor $A\mapsto Z^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times}))$. This section is the technical heart of the paper. In Section 5 we relate the functor $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ to the functor $Z^{1}(\Gamma_{1},G^{0}(-))$. In Section 6 we introduce Lafforgue’s $G$-pseudocharacters and study the functor $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$. In Section 7 we transfer the results about abstract representations and $G$-pseudocharacters of $\Gamma_{1}$ to continuous representations and continuous $G$-pseudocharacters of its profinite completion $\widehat{\Gamma}_{1}$ and prove versions of Theorems 1.3, 1.4, 1.5 with $\Gamma_{F}$ replaced by $\widehat{\Gamma}_{1}$. In Section 8 we compute $\mathbb{Z}_{p}$-ranks of certain pro-$p$ completions appearing naturally in the arithmetic setting. We show in Lemma 8.7 that the pro-$p$ completion of $(E^{\times}\otimes M)^{\Delta}$ is isomorphic to $(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}$. In Section 9 we use the results proved in the previous sections to prove Theorems 1.1, 1.3, 1.4, 1.5.
1.5. Notation
[…]
Let $\mathfrak{A}_{\mathcal{O}}$ be the category of local artinian $\mathcal{O}$-algebras with residue field $k$. We let $\widehat{\mathfrak{A}}_{\mathcal{O}}$ be the category of pro-objects of $\mathfrak{A}_{\mathcal{O}}$. Concretely, one may identify $\widehat{\mathfrak{A}}_{\mathcal{O}}$ with the category of pseudo-compact local $\mathcal{O}$-algebras with residue field $k$. Let $\mathfrak{A}_{k}$ be the full subcategory of $\mathfrak{A}_{\mathcal{O}}$ consisting of objects killed by $\varpi$. Then $\widehat{\mathfrak{A}}_{k}$ is the full subcategory of $\widehat{\mathfrak{A}}_{\mathcal{O}}$ consisting of objects killed by $\varpi$, which can be identified with the category of local pseudo-compact $k$-algebras with residue field $k$.
The groups denoted by $\Gamma_{1}$ and $\Gamma_{2}$ are abstract groups with no topology. We will denote by $\widehat{\Gamma}_{1}$ the profinite completion of $\Gamma_{1}$. If $\mathcal{A}$ is an abelian group then we will denote its pro-$p$ completion by $\mathcal{A}^{\wedge,p}$.
We denote by $\operatorname{Set}$, $\operatorname{Ab}$, $\operatorname{CRing}$ the categories of sets, abelian groups and commutative rings, respectively.
If $X$ is a scheme the we denote the ring of its global sections by $\mathcal{O}(X)$. If $\Gamma$ is a group then we denote by $\mathcal{O}[\Gamma]$ its group algebra over a ring $\mathcal{O}$.
1.6. Acknowledgements
VP would like to thank Timo Richarz for a helpful discussion concerning the Langlands correspondence for tori. The authors would like to thank Toby Gee and James Newton for their comments on a draft version of this paper.
Parts of the paper were written during the research stay of VP at the Hausdorff Research Institute for Mathematics in Bonn for the Trimester Program The Arithmetic of the Langlands Program. VP would like to thank the organisers Frank Calegari, Ana Caraiani, Laurent Fargues and Peter Scholze for the invitation and a stimulating research environment. The research stay of VP was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2047/1 – 390685813.
The research of JQ was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number 517234220.
2. Generalised tori
Definition 2.1.
Let $S$ be a scheme. A generalised torus is a smooth affine $S$-group scheme $G$, such that the geometric fibres of $G^{0}$ are tori and $G/G^{0}\to S$ is finite.
Remark 2.2.
It follows from [5, Proposition 3.1.3], that for a generalised torus $G$ the quotient $G/G^{0}$ is étale. Sean Cotner has shown in [6] that if $S$ is locally noetherian then for a smooth affine $S$-group scheme $G$ the identity component $G^{0}$ has reductive geometric fibers and $G/G^{0}$ is finite étale if and only if $G$ is geometrically reductive in the sense of [1, Definition 9.1.1]. In particular, this holds over $S=\operatorname{Spec}\mathcal{O}$.
Remark 2.3.
When $G$ is a generalised torus over $\mathcal{O}$, then $G^{0}$ is linearly reductive, i.e. taking $G^{0}$-invariants is an exact functor on the category of $\mathcal{O}(G^{0})$-comodules. It follows, that if $A$ is a commutative $\mathcal{O}$-algebra with trivial $G^{0}$-action, $M$ is an $A$-module with $G^{0}$-action and $N$ is an $A$-module with trivial $G^{0}$-action, then
${(M\otimes_{A}N)^{G^{0}}=M^{G^{0}}\otimes_{A}N}.$
If $M$ is an abelian group then we denote by $\mathfrak{D}(M):=\operatorname{Spec}(\mathcal{O}[M])$ the diagonalizable group scheme, where the Hopf algebra structure on $\mathcal{O}[M]$ is given by addition in $M$ and inversion. For any $\mathcal{O}$-algebra $A$, we have
$\mathfrak{D}(M)(A)=\operatorname{Hom}(M,A^{\times}).$
If $M$ is a finitely generated free abelian group then $\mathfrak{D}(M)$ is a split torus.
3. Shapiro’s lemma
[…]
Proposition 3.1.
Let $\Phi:\Gamma_{1}\rightarrow\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}$ be a $1$-cocycle, then the function $\varphi:\Gamma_{2}\rightarrow V$, given by
$\varphi(k)=[\Phi(k)](1)$
is a $1$-cocycle. Moreover, the above map induces an isomorphism in cohomology
$H^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V})\cong H^{1}(\Gamma_{2},V).$
Proof.
[9, Theorem 3.7]. ∎
We will construct an explicit inverse of the map above following [9]. For each right coset $c$ of $\Gamma_{2}$ in $\Gamma_{1}$ we fix a coset representative $\overline{c}$, so that the representative of the trivial coset is $1$. In particular,
$\Gamma_{1}=\bigcup_{c}\Gamma_{2}\overline{c}=\bigcup_{c}\overline{c}^{-1}\Gamma_{2}.$
Since $cg=\Gamma_{2}\overline{cg}$ for every $g\in\Gamma_{1}$, we have
$\overline{c}g\overline{cg}^{-1}\in\Gamma_{2},\quad\forall g\in\Gamma_{1}.$
Let $\mathcal{F}_{V}$ be the subgroup of $\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}$ of functions with support in $\Gamma_{2}$. By evaluating at $1$ we obtain a canonical isomorphism of $\Gamma_{2}$-representations between $\mathcal{F}_{V}$ and $V$; the inverse homomorphism is obtained by mapping $v$ to the function $g\mapsto gv$. If $\varphi\in Z^{1}(\Gamma_{2},V)$ then the isomorphism gives us a cocycle $f_{\varphi}\in Z^{1}(\Gamma_{2},\mathcal{F}_{V})$, where if $h\in\Gamma_{2}$ then $f_{\varphi}(h)\in\mathcal{\mathcal{F}}_{V}$ is the function given by
(7) $[f_{\varphi}(h)](k)=k\varphi(h),\quad\forall k\in\Gamma_{2}.$
Lemma 3.2.
Let $\varphi:\Gamma_{2}\rightarrow V$ be a $1$-cocycle, and let $f_{\varphi}:\Gamma_{2}\rightarrow\mathcal{F}_{V}$ be a $1$-cocycle corresponding to $\varphi$, via the canonical isomorphism $\mathcal{F}_{V}\cong V$ and let ${\Phi_{\varphi}:\Gamma_{1}\rightarrow\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}}$ be a function given by
$\Phi_{\varphi}(g)=\sum_{c}\overline{c}^{-1}f_{\varphi}(\overline{c}g\overline{cg}^{-1}).$
Then $\Phi_{\varphi}$ is a $1$-cocycle, such that
$[\Phi_{\varphi}(k)](1)=\varphi(k),\quad\forall k\in\Gamma_{2}.$
Proof.
We have to show that
$\Phi_{\varphi}(gh)=\Phi_{\varphi}(g)+g\Phi_{\varphi}(h),\quad\forall g,h\in\Gamma_{1}.$
It is enough to show that the equality holds once we evaluate both sides at $k\overline{c}$, where $\overline{c}$ is a representative of the coset $c$ and $k\in\Gamma_{2}$. If $f\in\mathcal{F}_{V}$ then $\operatorname{Supp}\overline{c}^{-1}f\subseteq c$, hence
$[\Phi_{\varphi}(gh)](k\overline{c})=[f_{\varphi}(\overline{c}gh\overline{cgh}^{-1})](k),\quad[\Phi_{\varphi}(g)](k\overline{c})=[f_{\varphi}(\overline{c}g\overline{cg}^{-1})](k),$
$\begin{split}[g\Phi_{\varphi}(h)](k\overline{c})=[\Phi_{\varphi}(h)](k\overline{c}g)=&[f_{\varphi}(\overline{cg}h\overline{cgh}^{-1})](k\overline{c}g\overline{cg}^{-1})\\ =&[\overline{c}g\overline{cg}^{-1}f_{\varphi}(\overline{cg}h\overline{cgh}^{-1})](k).\end{split}$
Since $f_{\varphi}$ is a $1$-cocycle we have
$\overline{c}g\overline{cg}^{-1}f_{\varphi}(\overline{cg}h\overline{cgh}^{-1})+f_{\varphi}(\overline{c}g\overline{cg}^{-1})=f_{\varphi}(\overline{c}g\overline{cg}^{-1}\overline{cg}h\overline{cgh}^{-1})=f_{\varphi}(\overline{c}gh\overline{cgh}^{-1})$
and hence $\Phi_{\varphi}$ is a $1$-cocycle. Since the representative of the trivial coset was chosen to be $1$, we have
$[\Phi_{\varphi}(k)](1)=[f_{\varphi}(k)](1)=\varphi(k),\quad\forall k\in\Gamma_{2}.$
Remark 3.3.
The map $\varphi\mapsto\Phi_{\varphi}$ depends on the choice of the coset representatives $\bar{c}$ and so it is not canonical. However once these representatives have been fixed it is immediate from the formulas that the map is functorial in $V$. So that if $\alpha:V\rightarrow W$ is a $\Gamma_{2}$-equivariant homomorphism of abelian groups equipped with $\Gamma_{2}$-action and $\psi\in Z^{1}(\Gamma_{2},W)$ is the image of $\varphi$ under the map $Z^{1}(\Gamma_{2},V)\rightarrow Z^{1}(\Gamma_{2},W)$ induced by $\alpha$ then $\Phi_{\varphi}$ maps to $\Phi_{\psi}\in Z^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{W})$.
4. The space of $1$-cocycles
We shall later identify our space of deformations with a space of $1$-cocycles. In this section we study the space of $1$-cocycles in a more abstract situation. We fix the following notation:
  • $\Gamma_{1}$ is an abstract group.
  • $M$ is an abelian group equipped with a linear left $\Gamma_{1}$-action with kernel $\Gamma_{2}$ of finite index in $\Gamma_{1}$.
[…]
We start by constructing an abelian group $\mathcal{E}$, which represents the functor
$\operatorname{Ab}\to\operatorname{Ab},~{}V\mapsto Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V).$
It turns out, that $\mathcal{E}$ carries an action of $\Delta$, such that $(\mathcal{E}\otimes M)_{\Delta}$ represents the functor
$\operatorname{Ab}\rightarrow\operatorname{Ab},~{}V\mapsto Z^{1}(\Gamma_{1},\operatorname{Hom}(M,V)).$
Our strategy is to realize $\operatorname{Hom}(M,V)$ as $\Delta$-invariants in $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)$, and then use the explicit version of Shapiro’s lemma of Section 3 to compute
$Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)).$
At the end of this section we study the functors $\operatorname{CRing}\rightarrow\operatorname{Ab}$ defined by
$A\mapsto Z^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times})),\quad A\mapsto H^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times})).$
Let $V$ be an abelian group equipped with a trivial $\Gamma_{2}$-action. We may identify $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ with the space of functions $f:\Delta\rightarrow V$ and $[g\cdot f](c)=f(cg)$ for all $g\in\Gamma_{1}$ and $c\in\Delta$. We also have an action of $\Delta$ on $\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}V$, which commutes with the action of $\Gamma_{1}$ and is given by $[d\cdot f](c)=f(d^{-1}c)$ for all $d\in\Delta$. This induces an action of $\Delta$ on $Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)$, which is given explicitly by
(8) $[[d\ast\Phi](g)](c)=[\Phi(g)](d^{-1}c).$
Since the action of $\Gamma_{2}$ on $V$ is trivial we have
(9) $Z^{1}(\Gamma_{2},V)=H^{1}(\Gamma_{2},V)=\operatorname{Hom}(\Gamma_{2},V),$
where $\operatorname{Hom}$ stands for group homomorphisms. For each $c\in\Delta$ we fix a coset representative $\overline{c}\in\Gamma_{1}$ as in the previous section.
Corollary 4.1.
Let $\varphi\in\operatorname{Hom}(\Gamma_{2},V)$. Then the $1$-cocycle $\Phi_{\varphi}$ constructed in 3.2 is given by
$\Phi_{\varphi}(g)=\sum_{c\in\Delta}\varphi(\overline{c}g\overline{cg}^{-1})\mathbf{1}_{c},$
where $\mathbf{1}_{c}$ is the indicator function for the coset $c$.
Proof.
Since the action of $\Gamma_{2}$ on $V$ is trivial by assumption it follows from (7) that $[f_{\varphi}(h)](k)=\varphi(h)$ for all $k,h\in\Gamma_{2}$. The assertion then follows from 3.2. ∎
3.1 and (9) give us an exact sequence of abelian groups:
$0\rightarrow B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\rightarrow Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\rightarrow\operatorname{Hom}(\Gamma_{2},V)\rightarrow 0$
and 3.2 gives us a section so that
(10) $Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\cong B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\oplus\operatorname{Hom}(\Gamma_{2},V)$
We want to understand the action of $\Delta$ on $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\oplus\operatorname{Hom}(\Gamma_{2},V)$ via this isomorphism. It follows from (8) that $[[d\ast\Phi](g)](cd)=[\Phi(g)](c)$. Hence
$[d\ast\Phi_{\varphi}](g)=\sum_{c\in\Delta}\varphi(\overline{c}g\overline{cg}^{-1})\mathbf{1}_{cd}.$
for $\varphi\in\operatorname{Hom}(\Gamma_{2},V)$. Since $\Gamma_{2}$ is normal in $\Gamma_{1}$, $ck=c$ for all $k\in\Gamma_{2}$ and all $c\in\Delta$. Hence, $\overline{ck}=\overline{c}$ and we conclude that
$[[d\ast\Phi_{\varphi}](k)](1)=\sum_{c}\varphi(\overline{c}k\overline{c}^{-1})\mathbf{1}_{cd}(1)=\varphi(\overline{d^{-1}}k\overline{d^{-1}}^{-1})=\varphi(\overline{d}^{-1}k\overline{d})=[d\cdot\varphi](k).$
We note that $\varphi((\overline{d}h)^{-1}k(\overline{d}h))=\varphi(\overline{d}^{-1}k\overline{d})$ for all $h,k\in\Gamma_{2}$, which justifies the last two equalities. Thus the difference $d\ast\Phi_{\varphi}-\Phi_{d\cdot\varphi}$ is a $1$-coboundary and we want to find a function $f:\Delta\rightarrow V$ depending on $d$ such that for all $g\in\Gamma_{1}$ we have
(11) $[d\ast\Phi_{\varphi}](g)-\Phi_{d\cdot\varphi}(g)=(g-1)f.$
After subtracting a constant function we may assume that $f(1)=0$. By evaluating both sides of (11) at $1$, we obtain:
(12) $\begin{split}f(g)&=f(g)-f(1)=[[d\ast\Phi_{\varphi}](g)](1)-[\Phi_{d\cdot\varphi}(g)](1)\\ &=\varphi(\overline{d^{-1}}g\overline{d^{-1}g}^{-1})-[d\cdot\varphi](g\overline{g}^{-1})\\ &=\varphi(\overline{d^{-1}}g\overline{d^{-1}g}^{-1})-\varphi(\overline{d^{-1}}g\overline{g}^{-1}\overline{d^{-1}}^{-1})\\ &=\varphi((\overline{d^{-1}}g\overline{g}^{-1}\overline{d^{-1}}^{-1})^{-1}\overline{d^{-1}}g\overline{d^{-1}g}^{-1})\\ &=\varphi(\overline{d^{-1}}\overline{g}\overline{d^{-1}g}^{-1}).\end{split}$
We note that in the above calculation $\overline{g}$ is the fixed coset representative of the coset $\Gamma_{2}g$. In particular, $\overline{g}\in\Gamma_{1}$ and not in $\Delta$. The function $f$ satisfies $f(kg)=f(g)$ for all $k\in\Gamma_{2}$, so we may consider as a function on $\Delta$. In this case the formula (12) applied with $g=\overline{c}$ gives us
(13) $f(c)=\varphi(\overline{d^{-1}}\overline{c}\overline{d^{-1}c}^{-1}),\quad\forall c\in\Delta.$
The function $f$ corresponds to a group homomorphism $\alpha_{f}:\mathbb{Z}[\Delta]\rightarrow V$, given by $\alpha_{f}(c)=f(c)$. Since $f(1)=0$ the restriction of $\alpha_{f}$ to the augmentation ideal $I_{\Delta}$ is given by
(14) $\alpha_{f}(c-1)=\varphi(\overline{d^{-1}}\overline{c}\overline{d^{-1}c}^{-1}),\quad\forall c\in\Delta.$
Since $V$ is abelian we have $\operatorname{Hom}(\Gamma_{2},V)=\operatorname{Hom}(\Gamma_{2}^{\mathrm{ab}},V)$, where $\Gamma_{2}^{\mathrm{ab}}$ is the maximal abelian quotient of $\Gamma_{2}$. For $d,c\in\Delta$ we define $\kappa(d,c)$ to be the image of $\overline{d}\overline{c}\overline{dc}^{-1}$ in $\Gamma_{2}^{\mathrm{ab}}$. Then $\kappa:\Delta\times\Delta\rightarrow\Gamma_{2}^{\mathrm{ab}}$ is the $2$-cocycle corresponding to the extension
$0\rightarrow\Gamma_{2}^{\mathrm{ab}}\rightarrow\Gamma_{1}/[\Gamma_{2},\Gamma_{2}]\rightarrow\Delta\rightarrow 0$
We define an action of $\Delta$ on $\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta}$ by letting
(15) $\begin{split}d\ast(g,c-1):=(dgd^{-1}\cdot\kappa(d,c),dc-d).\end{split}$
This action defines an action of $\Delta$ on $\operatorname{Hom}(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta},V)$ by
(16) $\begin{split}[d\ast(\varphi,\alpha)](g,c-1)&=(\varphi,\alpha)(d^{-1}\ast(g,c-1))\\ &=\varphi(d^{-1}gd)+\varphi(\kappa(d^{-1},c))+\alpha(d^{-1}c-d^{-1}).\end{split}$
We obtain an exact sequence of $\mathbb{Z}[\Delta]$-modules
(17) $\begin{split}0\rightarrow\Gamma_{2}^{\mathrm{ab}}\rightarrow\mathcal{E}\rightarrow I_{\Delta}\rightarrow 0.\end{split}$
by letting $\mathcal{E}=\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta}$ with the $\Delta$-action defined as above.
Remark 4.2.
Using the exact sequence $0\rightarrow I_{\Delta}\rightarrow\mathbb{Z}[\Delta]\rightarrow\mathbb{Z}\rightarrow 0$ of $\mathbb{Z}[\Delta]$-modules one obtains a canonical identification
$\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})\cong\operatorname{Ext}^{2}_{\mathbb{Z}[\Delta]}(\mathbb{Z},\Gamma_{2}^{\mathrm{ab}})\cong H^{2}(\Delta,\Gamma_{2}^{\mathrm{ab}}),$
see [3, Chapter XIV, §4, Remark]. The extension class of (17) gives an element in $\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})$. One can show, that the image of this class in $H^{2}(\Delta,\Gamma_{2}^{\mathrm{ab}})$ is equal to the class of $\kappa$.
[…]
(18) $\operatorname{Hom}(I_{\Delta},V)\rightarrow B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V),\quad\alpha\mapsto b_{\alpha}$
is an isomorphism. We now record the consequence of our calculations:
Proposition 4.3.
If $V$ is an abelian group with the trivial $\Gamma_{2}$-action then sending $(\varphi,\alpha)$ to $\Phi_{\varphi}+b_{\alpha}$ induces an isomorphism
$\operatorname{Hom}(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta},V)\overset{\cong}{\longrightarrow}Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V),$
which is $\Delta$-equivariant for the actions defined in (16) and (8).
Proof.
It follows from (18) and (10) that the map is an isomorphism of abelian groups. We have to check that it is $\Delta$-equivariant. We have
$\begin{split}[[d\ast b_{\alpha}](g)](c)&=[b_{\alpha}(g)](d^{-1}c)=f_{\alpha}(d^{-1}cg)-f_{\alpha}(d^{-1}c)\\ &=\alpha(d^{-1}cg-1)-\alpha(d^{-1}c-1)=\alpha(d^{-1}cg-d^{-1}c),\end{split}$
$\begin{split}[b_{d\ast\alpha}(g)](c)&=f_{d\ast\alpha}(cg)-f_{d\ast\alpha}(c)=[d\ast\alpha](cg-1)-[d\ast\alpha](c-1)\\ &=\alpha(d^{-1}cg-d^{-1})-\alpha(d^{-1}c-d^{-1})=\alpha(d^{-1}cg-d^{-1}c).\end{split}$
Hence, $d\ast b_{\alpha}=b_{d\ast\alpha}$. It follows from (11), (13) and (14) that $d\ast\Phi_{\varphi}=\Phi_{d\cdot\varphi}+b_{\beta}$, where $\beta(c-1)=\varphi(\kappa(d^{-1},c))$ for all $c\in\Delta$. On the other hand, it follows from (16) that $d\ast(\varphi,0)=(d\cdot\varphi,\beta)$. Hence, $d\ast(\varphi,0)$ is mapped to $\Phi_{d\cdot\varphi}+b_{\beta}=d\ast\Phi_{\varphi}$. ∎
Recall, that $M$ is an abelian group with an action of $\Delta$, so $\Gamma_{2}$ acts trivially on $\operatorname{Hom}(M,V)$. We have natural isomorphisms of $\Gamma_{1}$-representations
(19) $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)\cong\operatorname{Hom}(\mathbb{Z}[\Delta],\operatorname{Hom}(M,V))\cong\operatorname{Hom}(\mathbb{Z}[\Delta]\otimes M,V),$
where the action of $\Gamma_{1}$ on the last term is induced its action on $\mathbb{Z}[\Delta]\otimes M$, which is given by $g\cdot(c\otimes m)=cg^{-1}\otimes m$. The diagonal action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$ is given by $d\cdot(c\otimes m)=dc\otimes dm$, commutes with the action of $\Gamma_{1}$ and induces an action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)$, which commutes with the action of $\Gamma_{1}$. If $M=\mathbb{Z}$ with the trivial $\Delta$-action then the construction recovers the action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ considered earlier.
Lemma 4.4.
$(\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))^{\Delta}$ as $\Gamma_{1}$-representation is naturally isomorphic to $\operatorname{Hom}(M,V)$ with $\Gamma_{1}$ acting on it via $\Delta$.
Proof.
The map $\mathbb{Z}[\Delta]\otimes M\rightarrow\mathbb{Z}[\Delta]\otimes M$ , $c\otimes m\mapsto c\otimes c^{-1}m$ is an isomorphism of $\Delta\times\Gamma_{1}$-representations, where on the source the action is the one considered above, and on the target $\Delta$ acts by $d\ast(c\otimes m)=dc\otimes m$ and $\Gamma_{1}$ acts by $g\ast(c\otimes m)=cg^{-1}\otimes gm$. It is immediate by considering the $\ast$-action that $(\mathbb{Z}[\Delta]\otimes M)_{\Delta}\cong M$ and $\Gamma_{1}$ acts on $M$ via $\Delta$. Using (19) we can then translate this statement to a statement about invariants. ∎
Proposition 4.5.
Let $\mathcal{E}$ be the $\mathbb{Z}[\Delta]$-module constructed above. Then for all abelian groups $V$, we have an isomorphism of abelian groups functorial in $V$:
(20) $\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},V)\cong Z^{1}(\Gamma_{1},\operatorname{Hom}(M,V)),$
where the action of $\Delta$ on $\mathcal{E}\otimes M$ is diagonal and $\Gamma_{1}$ acts on $\operatorname{Hom}(M,V)$ via $\Delta$.
Proof.
4.3 gives us $\Delta$-equivariant isomorphisms:
$\operatorname{Hom}(\mathcal{E}\otimes M,V)\cong\operatorname{Hom}(\mathcal{E},\operatorname{Hom}(M,V))\cong Z^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)).$
The assertion follows after taking $\Delta$-invariants and applying 4.4. ∎
The exact sequence $0\rightarrow I_{\Delta}\rightarrow\mathbb{Z}[\Delta]\rightarrow\mathbb{Z}\rightarrow 0$ remains exact after tensoring with $M$. Taking $\Delta$-coinvariants yields an exact sequence
$0\rightarrow H_{1}(\Delta,M)\rightarrow(I_{\Delta}\otimes M)_{\Delta}\rightarrow M\rightarrow M_{\Delta}\rightarrow 0$
This gives us a surjection $(I_{\Delta}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$. By composing this map with the surjection $(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow(I_{\Delta}\otimes M)_{\Delta}$ induced by (17) we obtain a surjection $(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$. This in turn yields an injection
$\operatorname{Hom}(I_{\Delta}M,V)\hookrightarrow\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},V).$
Lemma 4.6.
The isomorphism (20) in 4.5 identifies the space of $1$-coboundaries $B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ with a subgroup of $\operatorname{Hom}(I_{\Delta}M,V)$. Moreover, it induces an isomorphism between the two groups if $V$ is divisible.
Proof.
The isomorphism $(\mathbb{Z}[\Delta]\otimes M)_{\Delta}\cong M$ for the diagonal action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$ is realised by the map $c\otimes m\mapsto c^{-1}m$. The image of $I_{\Delta}\otimes M$ under this map is equal to $I_{\Delta}M$. Thus the isomorphism
$\vartheta:\operatorname{Hom}(M,V)\overset{\cong}{\longrightarrow}(\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))^{\Delta}$
in 4.4 is given explicitly by
$[\vartheta(\alpha)(c)](m)=\alpha(c^{-1}m),\quad\forall c\in\Delta,\quad\forall m\in M.$
Let $b\in B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ be a boundary $b(g)=(g-1)\alpha$. Its image in $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))$ is the boundary $b^{\prime}$ given by $b^{\prime}(g)=(g-1)\vartheta(\alpha)$. The constant function $f^{\prime}(c)=\alpha$ for all $c\in\Delta$ is $\Gamma_{1}$-invariant thus
$b^{\prime}(g)=(g-1)(\vartheta(\alpha)-f^{\prime}),\quad\forall g\in\Gamma_{1}.$
Since $(\vartheta(\alpha)-f^{\prime})(1)=\alpha-\alpha=0$, by 4.3 $b^{\prime}=b_{\beta}$, where $\beta:I_{\Delta}\rightarrow\operatorname{Hom}(M,V)$ is given by
$[\beta(c-1)](m)=[(\vartheta(\alpha)-f^{\prime})(c)](m)=\alpha(c^{-1}m)-\alpha(m)=\alpha(c^{-1}m-m).$
We conclude that the image of $B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ under the isomorphism in 4.5 is contained in $\operatorname{Hom}(I_{\Delta}M,V)$.
Conversely, if we start with a homomorphism $\beta^{\prime}:I_{\Delta}M\rightarrow V$ then by pulling it back under the surjection $I_{\Delta}\otimes M\twoheadrightarrow I_{\Delta}M$ we obtain a homomorphism $\beta:I_{\Delta}\otimes M\rightarrow V$ given by $\beta((c-1)\otimes m)=\beta^{\prime}(c^{-1}m-m)$. The corresponding coboundary $b_{\beta}\in B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))$ is given by $b_{\beta}(g)=(g-1)f_{\beta}$, where
$[f_{\beta}(c)](m)=\beta((c-1)\otimes m)=\beta^{\prime}(c^{-1}m-m).$
If $V$ is divisible then $V$ is injective in $\operatorname{Ab}$ and the exact sequence
$0\rightarrow I_{\Delta}M\rightarrow M\rightarrow M_{\Delta}\rightarrow 0$
gives rise to an exact sequence
$0\rightarrow\operatorname{Hom}(M_{\Delta},V)\rightarrow\operatorname{Hom}(M,V)\rightarrow\operatorname{Hom}(I_{\Delta}M,V)\rightarrow 0$
and we conclude that there is $\alpha\in\operatorname{Hom}(M,V)$ mapping to $\beta^{\prime}$, so that
$\beta^{\prime}(c^{-1}m-m)=\alpha(c^{-1}m)-\alpha(m).$
Let $f^{\prime}$ be the constant function $f^{\prime}(c)=\alpha$ for all $c\in\Delta$. Then the boundary $g\mapsto(g-1)(f_{\beta}+f^{\prime})$ is equal to $b_{\beta}$ and $f_{\beta}+f^{\prime}=\vartheta(\alpha)$. ∎
By taking $V=(\mathcal{E}\otimes M)_{\Delta}$ in 4.5 we obtain a obtain a natural cocycle
$\Phi_{\mathrm{nat}}\in Z^{1}(\Gamma_{1},\operatorname{Hom}(M,(\mathcal{E}\otimes M)_{\Delta})),$
which corresponds to the identity in $\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},(\mathcal{E}\otimes M)_{\Delta})$.
Lemma 4.7.
If $V$ is any abelian group then for all $\Gamma_{1}$-modules $N$ we have a canonical map
(21) $H^{1}(\Gamma_{1},\operatorname{Hom}(N,V))\rightarrow\operatorname{Hom}(H_{1}(\Gamma_{1},N),V),$
which is an isomorphism if $V$ is divisible.
Proof.
If $(C_{\bullet},d_{\bullet})$ is a complex of abelian groups then we let $Z_{n}=\ker(d_{n}:C_{n}\rightarrow C_{n-1})$ and $Z^{n}=\ker(d_{n+1}^{\ast}:\operatorname{Hom}(C_{n},V)\rightarrow\operatorname{Hom}(C_{n+1},V))$. The evaluation pairing $Z_{n}\times Z^{n}\rightarrow V$ induces a bilinear map $H_{n}(C_{\bullet})\times H^{n}(\operatorname{Hom}(C_{\bullet},V))\rightarrow V$, which induces a canonical map $H^{n}(\operatorname{Hom}(C_{\bullet},V))\rightarrow\operatorname{Hom}(H_{n}(C_{\bullet}),V)$. If $V$ is divisible then $V$ is injective in $\operatorname{Ab}$ and the map is an isomorphism.
[…]
Applying 4.7 with $N=M$ and $V=(\mathcal{E}\otimes M)_{\Delta}$ we obtain a homomorphism $\varphi_{\mathrm{nat}}:H_{1}(\Gamma_{1},M)\rightarrow(\mathcal{E}\otimes M)_{\Delta}$ corresponding to the cohomology class $[\Phi_{\mathrm{nat}}]$.
Lemma 4.8.
For all abelian groups $V$ the composition
[…]
is given by $\psi\mapsto\psi\circ\varphi_{\mathrm{nat}}$.
Proof.
This follows from the fact that all our constructions are functorial in $V$. If we denote the four functors appearing in (22) with $A$, $B$, $C$, $D$ then for a homomorphism of abelian groups $\psi:W\rightarrow V$ we obtain a diagram:
${A(W)}$${B(W)}$${C(W)}$${D(W)}$${A(V)}$${B(V)}$${C(V)}$${D(V)}$$\scriptstyle{\psi\circ}$$\scriptstyle{\psi\circ}$$\scriptstyle{\psi\circ}$$\scriptstyle{\psi\circ}$
with commutative squares. If we take $W=(\mathcal{E}\otimes M)_{\Delta}$ then the identity map in $A((\mathcal{E}\otimes M)_{\Delta})$ maps to $\varphi_{\mathrm{nat}}$ in $D((\mathcal{E}\otimes M)_{\Delta})$ by construction and hence to $\psi\circ\varphi_{\mathrm{nat}}$ in $D(V)$. Since the identity maps to $\psi$ in $A(V)$ we obtain the assertion. ∎
Proposition 4.9.
There is an exact sequence of abelian groups
[…]
which is functorial in $M$ and for all abelian groups $V$ the diagram
[…]
commutes.
Proof.
Lemmas 4.6 and 4.8 give us the commutative diagram above. Moreover, if $V$ is divisible then the top row is exact, the last arrow in the top row is surjective and the first vertical arrow is an isomorphism. We deduce that for all divisible $V$ the maps $q:(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$ and $\varphi_{\mathrm{nat}}:H_{1}(\Gamma,M)\rightarrow(\mathcal{E}\otimes M)_{\Delta}$ induce an exact sequence
(24) $0\rightarrow\operatorname{Hom}(I_{\Delta}M,V)\rightarrow\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},V)\rightarrow\operatorname{Hom}(H_{1}(\Gamma,M),V)\rightarrow 0$
Since divisible groups are precisely injective objects in $\operatorname{Ab}$, which has enough injectives, an abelian group $A$ is zero if and only if $\operatorname{Hom}(A,V)=0$ for all divisible groups $V$. Using this and the exactness of (24) we obtain that $\mathrm{Im}(q\circ\varphi_{\mathrm{nat}})=0$, thus (23) is a complex and a further application of the same argument shows that (23) is exact. ∎
Corollary 4.10.
The functor $\operatorname{CRing}\rightarrow\operatorname{Ab}$ given by
$A\mapsto Z^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times}))$
is represented by the group algebra $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]$.
Proof.
If $W$ is an abelian group then we may identify $W$ with a subgroup of units of the group ring $\mathbb{Z}[W]$. The map
$\operatorname{Hom}_{\operatorname{CRing}}(\mathbb{Z}[W],A)\rightarrow\operatorname{Hom}(W,A^{\times}),\quad\psi\mapsto\psi|_{W}$
is an isomorphism. Applying this observation to $W=(\mathcal{E}\otimes M)_{\Delta}$ and using 4.5 yields the assertion. ∎
Corollary 4.11.
If $M$ is a free $\mathbb{Z}$-module of finite rank then
(25) $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]\cong\mathbb{Z}[H_{1}(\Gamma,M)][t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
where $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
Proof.
Since $I_{\Delta}M$ is a submodule of $M$, which is free of finite rank over $\mathbb{Z}$, we have an isomorphism $I_{\Delta}M\cong\mathbb{Z}^{s}$, where $s$ is as above. By choosing a section to the surjection $(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$ in (23) we obtain an isomorphism $(\mathcal{E}\otimes M)_{\Delta}\cong H_{1}(\Gamma_{1},M)\times\mathbb{Z}^{s}$ and this implies the assertion. ∎
Proposition 4.12.
The map (21) induces a map of presheaves $\operatorname{CRing}\to\operatorname{Set}$
(26) $H^{1}(\Gamma_{1},\operatorname{Hom}(M,(-)^{\times}))\to\operatorname{Spec}(\mathbb{Z}[H_{1}(\Gamma_{1},M)])$
which exhibits the right hand side as the fpqc sheafification of the left hand side.
Proof.
The right hand side is an fpqc sheaf, since it is representable [14, Tag 023Q]. So it suffices to find for every ring $A$ and every homomorphism ${f:\mathbb{Z}[H_{1}(\Gamma_{1},M)]\to A}$ an fpqc cover of $A$ and a descent datum in the left hand side, which maps to a descent datum in the right hand side, which descends to $f$. By [16, Lemma 4.1.1] for every ring $A$ there is a faithfully flat map $A\to B$, such that $B^{\times}$ is divisible. By 4.7 the map
$H^{1}(\Gamma_{1},\operatorname{Hom}(M,B^{\times}))\rightarrow\operatorname{Hom}_{\operatorname{CRing}}(\mathbb{Z}[H_{1}(\Gamma_{1},M)],B)$
is bijective. So the canonical descent datum associated with $f$ and the map $A\to B$ comes from a descent datum in the right hand side. ∎
5. Admissible representations
In this section, let $\mathcal{O}$ be an arbitrary commutative ring, such that $\operatorname{Spec}\mathcal{O}$ is connected. Let $G$ be a generalised torus over $\mathcal{O}$, such that $G^{0}$ is split and such that $G/G^{0}$ is constant. Let $\Pi:G\to G/G^{0}$ be the projection map. We define the character lattice of $G^{0}$ by
[…]
Let $\Gamma_{1}$ be an abstract group and let $\pi:\Gamma_{1}\twoheadrightarrow\Delta$ be a surjective homomorphism with kernel $\Gamma_{2}$. So we are in the situation of Section 4.
Definition 5.1.
We say that a representation $\rho:\Gamma_{1}\rightarrow G(A)$ is admissible if $\Pi\circ\rho:\Gamma_{1}\to\underline{\Delta}(A)$, sends $\gamma\in\Gamma_{1}$ to the image of $\pi(\gamma)\in\Delta$ in $\underline{\Delta}(A)$.
This terminology is motivated by admissible Galois representations into $L$-groups appearing in the Langlands correspondence.
Let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ be the set of admissible representations $\rho:\Gamma_{1}\rightarrow G(A)$. The group $\mathfrak{D}(M)(A)$ acts on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ by conjugation. This defines a scheme-theoretic action
(28) $\mathfrak{D}(M)\times\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}\rightarrow\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}.$
Proposition 5.2.
Assume, there is a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. Then
(29) $Z^{1}(\Gamma_{1},\mathfrak{D}(M)(A))\to\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A),\quad\Phi\mapsto[\gamma\mapsto\Phi(\gamma)\rho_{0}(\gamma)]$
is a bijection, which is natural in $A\in\mathcal{O}\text{-}\mathrm{alg}$. In particular, $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ is representable by an $\mathcal{O}$-algebra and (29) induces a natural isomorphism
(30) $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\xrightarrow{\cong}\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$
Moreover, (29) is $\mathfrak{D}(M)(A)$-equivariant and induces a natural bijection between the set of $\mathfrak{D}(M)(A)$-orbits in $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ and $H^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$. In particular, (30) is $\mathfrak{D}(M)$-equivariant.
Proof.
For $\Phi\in Z^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$, we have
(31) $\begin{split}\Phi(\gamma_{1}\gamma_{2})\rho_{0}(\gamma_{1}\gamma_{2})&=\Phi(\gamma_{1}){}^{\gamma_{1}}\Phi(\gamma_{2})\rho_{0}(\gamma_{1})\rho_{0}(\gamma_{2})\\ &=\Phi(\gamma_{1})\rho_{0}(\gamma_{1})\Phi(\gamma_{2})\rho_{0}(\gamma_{2}),\end{split}$
[…]
The $\mathfrak{D}(M)(A)$-conjugacy classes are the $B^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$-orbits: if we write an admissible homomorphism $\rho:\Gamma_{1}\to G(A)$ as above as $\rho(\gamma)=\Phi(\gamma)\rho_{0}(\gamma)$ and $g\in\mathfrak{D}(M)(A)$, then
$\begin{split}g\rho(\gamma)g^{-1}&=g\Phi(\gamma)({}^{\gamma}g^{-1})\rho_{0}(\gamma)\\ &=\beta_{g}(\gamma)\Phi(\gamma)\rho_{0}(\gamma),\end{split}$
where $\beta_{g}\in B^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$ is the coboundary $\beta_{g}(\gamma)=g({}^{\gamma}g^{-1})$. It follows that the $\mathfrak{D}(M)(A)$-orbits are the classes in $H^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$. ∎
Remark 5.3.
If $G=G^{0}\rtimes\underline{\Delta}$, then the composition of $\pi$ with the natural map $\Delta\to\underline{\Delta}(\mathcal{O})\to G^{0}(\mathcal{O})\rtimes\underline{\Delta}(\mathcal{O})=G(\mathcal{O})$ is a canonical choice for a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. In general, the isomorphisms (29) and (30) depend on $\rho_{0}$.
Proposition 5.4.
Assume, there is a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. Under the isomorphism (29) the action (28) corresponds to the ring homomorphism
$\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]\rightarrow\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]\otimes\mathcal{O}[M],$
which sends $x\in(\mathcal{E}\otimes M)_{\Delta}$ to $x\otimes q(x)$, where $q:(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$ is the map of (23).
Proof.
We have seen in the proof of 5.2, that the action of $\mathfrak{D}(M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ corresponds under the isomorphism (29) to the action via the boundary map
$\mathfrak{D}(M)(A)\to Z^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$
[…]
Corollary 5.5.
The action of $\mathfrak{D}(M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ factors through the action of $\mathfrak{D}(I_{\Delta}M)$, which acts freely.
Proof.
[…]
Corollary 5.6.
Assume that there is a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. The isomorphisms (26) and (29) induce an isomorphism
(32) $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}/\mathfrak{D}(M)\xrightarrow{\cong}\operatorname{Spec}(\mathcal{O}[H_{1}(\Gamma_{1},M)]),$
where $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}/\mathfrak{D}(M)$ denotes the fpqc sheafification of the presheaf quotient. Moreover this quotient is a GIT quotient. In particular, (32) induces isomorphisms
(33) $\mathcal{O}[H_{1}(\Gamma_{1},M)]\xrightarrow{\cong}\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})^{G^{0}}\xrightarrow{\cong}\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]^{G^{0}}$
and the composition of the maps in (33) is induced by $\varphi_{\mathrm{nat}}$.
Proof.
The isomorphism (32) follows directly from 4.12 and 5.2. To see that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}/\mathfrak{D}(M)$ is also a GIT quotient, we observe that every $\mathfrak{D}(M)$-equivariant map from $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ to an affine scheme equipped with the trivial $\mathfrak{D}(M)$-action factors through $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}/\mathfrak{D}(M)$ by the universal property of the fpqc sheaf quotient. Since $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}/\mathfrak{D}(M)$ is affine it is indeed the GIT quotient and this gives the first isomorphism in (33). The second isomorphism in (33) comes from (30).
Let $B\in\mathcal{O}\text{-}\mathrm{alg}$. After removing the $G^{0}$-invariants in (33) we get two maps
$\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},B^{\times})\xrightarrow{\cong}\operatorname{Rep}_{G,\pi}^{\Gamma_{1}}(B)\to\operatorname{Hom}(H_{1}(\Gamma_{1},M),B^{\times})$
The first map comes from (20) and (29). The second map comes from (21) and (29). Hence the composition is (22) with $V=B^{\times}$, so the last assertion follows from 4.8. ∎
Lemma 5.7.
Let $\tau:G\hookrightarrow\mathbb{A}^{n}$ be a closed immersion of $\mathcal{O}$-group schemes and let $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$. Assume that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. Then $\tau(\rho(\Gamma_{1}))$ is contained in a finitely generated $\mathcal{O}[H_{1}(\Gamma_{1},M)]$-submodule of $A^{n}=\mathbb{A}^{n}(A)$.
Proof.
Proposition 5.2 implies that it is enough to prove the statement, when $A=\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ and $\rho=\Phi_{\mathrm{nat}}\rho_{0}$.
Since $\Gamma_{2}$ is of finite index in $\Gamma_{1}$ it is enough to show that $\tau(\rho(\Gamma_{2}))$ is contained in a finitely generated $\mathcal{O}[H_{1}(\Gamma_{1},M)]$-submodule of $(\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}])^{n}$. Since $\Gamma_{2}$ acts trivially on $M$ we have a canonical isomorphism
$H_{1}(\Gamma_{2},M)\cong\Gamma_{2}^{\mathrm{ab}}\otimes M,$
and it follows from (23) that the image of $\mathcal{O}[\Gamma_{2}^{\mathrm{ab}}\otimes M]$ in $\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ is contained in $\mathcal{O}[H_{1}(\Gamma_{1},M)]$. We thus may assume that $G=\mathfrak{D}(M)$ and $\Gamma_{2}=\Gamma_{1}$. In this case, $\mathcal{O}[H_{1}(\Gamma,M)]=\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$, so there is nothing to prove. ∎
6. Lafforgue’s $G$-pseudocharacters
We keep the notation of Section 5. We now recall Lafforgue’s notion of $G$-pseudocharacter in the form of [12, Definition 3.1]. The difference to Lafforgue’s original definition [8, Section 11] is that we work over the base ring $\mathcal{O}$ and allow $G$ to be disconnected. The definition works for generalised reductive $\mathcal{O}$-group schemes as defined in [11], but this generality is not needed here.
Definition 6.1.
Let $A$ be a commutative $\mathcal{O}$-algebra. A $G$-pseudocharacter $\Theta$ of $\Gamma$ over $A$ is a sequence $(\Theta_{n})_{n\geq 1}$ of $\mathcal{O}$-algebra maps
$\Theta_{n}:\mathcal{O}[G^{n}]^{G^{0}}\to\mathrm{Map}(\Gamma^{n},A)$
for $n\geq 1$, satisfying the following conditions [^1] Here $G$ acts on $G^{n}$ by $g\cdot(g_{1},\dots,g_{n})=(gg_{1}g^{-1},\dots,gg_{n}g^{-1})$. This induces a rational action of $G$ on the affine coordinate ring $\mathcal{O}[G^{n}]$ of $G^{n}$. The submodule $\mathcal{O}[G^{n}]^{G^{0}}\subseteq\mathcal{O}[G^{n}]$ is defined as the rational invariant module of the $G^{0}$-representation $\mathcal{O}[G^{n}]$. It is an $\mathcal{O}$-subalgebra, since $G$ acts by $\mathcal{O}$-linear automorphisms.:
  1. (1)
    For each $n,m\geq 1$, each map $\zeta:\{1,\dots,m\}\to\{1,\dots,n\}$, $f\in\mathcal{O}[G^{m}]^{G^{0}}$ and $\gamma_{1},\dots,\gamma_{n}\in\Gamma$, we have
    $\Theta_{n}(f^{\zeta})(\gamma_{1},\dots,\gamma_{n})=\Theta_{m}(f)(\gamma_{\zeta(1)},\dots,\gamma_{\zeta(m)})$
    where $f^{\zeta}(g_{1},\dots,g_{n})=f(g_{\zeta(1)},\dots,g_{\zeta(m)})$.
  2. (2)
    For each $n\geq 1$, for each $\gamma_{1},\dots,\gamma_{n+1}\in\Gamma$ and each $f\in\mathcal{O}[G^{n}]^{G^{0}}$, we have
    $\Theta_{n+1}(\hat{f})(\gamma_{1},\dots,\gamma_{n+1})=\Theta_{n}(f)(\gamma_{1},\dots,\gamma_{n}\gamma_{n+1})$
    where $\hat{f}(g_{1},\dots,g_{n+1})=f(g_{1},\dots,g_{n}g_{n+1})$.
We denote the set of $G$-pseudocharacters of $\Gamma_{1}$ over $A$ by $\mathrm{PC}_{G}^{\Gamma_{1}}(A)$. The functor $A\mapsto\mathrm{PC}_{G}^{\Gamma_{1}}(A)$ is representable by an affine $\mathcal{O}$-scheme $\mathrm{PC}_{G}^{\Gamma_{1}}$ [12, Theorem 3.20]. When $\varphi:G\to H$ is a homomorphism of generalised tori over $\mathcal{O}$, the induced maps $\varphi^{*}_{n}:\mathcal{O}[H^{n}]^{H^{0}}\to\mathcal{O}[G^{n}]^{G^{0}}$ give rise to an $H$-pseudocharacter $(\Theta_{n}\circ\varphi^{*}_{n})_{n\geq 1}$. By analogy with the notation for representations we denote this $H$-pseudocharacter by $\varphi\circ\Theta$. So we also have an induced map $\mathrm{PC}_{G}^{\Gamma}(A)\to\mathrm{PC}_{H}^{\Gamma}(A)$. It is easy to verify that specialisation along $f:A\to B$ commutes with composition with $\varphi$, i.e. $(\varphi\circ\Theta)\otimes_{A}B=\varphi\circ(\Theta\otimes_{A}B)$.
Recall that for every homomorphism $\rho:\Gamma_{1}\to G(A)$, there is an associated $G$-pseudocharacter $\Theta_{\rho}\in\mathrm{PC}^{\Gamma_{1}}_{G}(A)$, which depends on $\rho$ only up to $G^{0}(A)$-conjugation. For $m\geq 1$, the homomorphism $(\Theta_{\rho})_{m}:\mathcal{O}(G^{m})^{G^{0}}\to\mathrm{Map}(\Gamma_{1}^{m},A)$ is defined by
[…]
There is a natural $G^{0}(A)$-equivariant map
(35) $\operatorname{Rep}^{\Gamma_{1}}_{G}(A)\to\mathrm{PC}^{\Gamma_{1}}_{G}(A),\quad\rho\mapsto\Theta_{\rho},$
which induces maps of $\mathcal{O}$-schemes
(36) $\operatorname{Rep}^{\Gamma_{1}}_{G}\to\mathrm{PC}^{\Gamma_{1}}_{G}.$
Since $G^{0}$-acts trivially on the target (36) factors as
(37) $\operatorname{Rep}^{\Gamma_{1}}_{G}\sslash G^{0}\to\mathrm{PC}^{\Gamma_{1}}_{G}.$
If $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$, then $\Theta_{\rho}$ is admissible in the sense, that the functorial image of $\Theta_{\rho}$ under the map $G\to\underline{\Delta}$ maps to the $\underline{\Delta}$-pseudocharacter attached to $\pi$ in $\mathrm{PC}^{\Gamma_{1}}_{\underline{\Delta}}(A)$. We denote the set of admissible $G$-pseudocharacters by $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$. The functor $A\mapsto\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$ is representable by an affine $\mathcal{O}$-scheme $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$. We have natural maps
(38) $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}\to\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$
(39) $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}\sslash G^{0}\to\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$
as above.
We will show, that (37) and (39) are isomorphisms. The result is specific to generalised tori [^2] We don’t expect (37) and (39) to be isomorphisms when $G$ is an arbitrary generalised reductive group in the sense of [12, Definition 2.3]. However, we don’t know of an example where (37) is not an isomorphism. and might be of independent interest.
Proposition 6.2.
The map
(40) $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})\to\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})^{G^{0}}$
corresponding to (37) is an isomorphism.
Proof.
As $G^{0}$ is linearly reductive the proof of [7, Proposition 2.11 (i)] applies, we recall the argument here. Let $\mathcal{O}(G^{\Gamma_{1}})$ be the $\mathcal{O}$-algebra which represents the functor
$\mathcal{O}\text{-}\mathrm{alg}\to\operatorname{Set},~{}A\mapsto\mathrm{Map}(\Gamma_{1},G(A)).$
We have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})\twoheadrightarrow\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})$ with kernel $J$. We also have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})^{G^{0}}\twoheadrightarrow\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})$ with kernel $I$. The structure of $J$ and $I$ is described explicitly in [7, Proposition 2.5]; note that there $G$ is assumed to be connected, but the description generalises easily to our situation. Namely $J$ is generated by the image of the $G^{0}$-equivariant maps
$\varphi_{\gamma,\delta}:\mathcal{O}(G\times G^{\Gamma_{1}})\to\mathcal{O}(G^{\Gamma_{1}})$
for all $\gamma,\delta\in\Gamma_{1}$, where
$\varphi_{\gamma,\delta}(f)((g_{\alpha})_{\alpha\in\Gamma_{1}})=f(g_{\gamma\delta},(g_{\alpha})_{\alpha\in\Gamma_{1}})-f(g_{\gamma}g_{\delta},(g_{\alpha})_{\alpha\in\Gamma_{1}}).$
The ideal $I$ is generated by the image of $\varphi_{\gamma,\delta}(\mathcal{O}(G\times G^{\Gamma_{1}})^{G^{0}})$ for all $\gamma,\delta\in\Gamma_{1}$.
Since taking $G^{0}$-invariants is exact, the natural map $\mathcal{O}(G^{\Gamma_{1}})^{G^{0}}/I=\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})\to\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})^{G^{0}}=\mathcal{O}(G^{\Gamma_{1}})^{G^{0}}/J^{G^{0}}$ is surjective. So it remains to show, that $J^{G^{0}}\subseteq I$. Let $h\in J^{G}$ and write $h=\sum_{i=1}^{n}\varphi_{\gamma_{i},\delta_{i}}(f_{i})$ where $\gamma_{i},\delta_{i}\in\Gamma_{1}$ and $f_{i}\in\mathcal{O}(G\times G^{\Gamma_{1}})$. Denote the Reynolds operator on $G^{0}$-modules by $E$, it commutes with the $G^{0}$-equivariant maps $\varphi_{\gamma_{i},\delta_{i}}$, so that we have
$h=E(h)=\sum_{i=1}^{n}\varphi_{\gamma_{i},\delta_{i}}(E(f_{i}))\in I.$
Lemma 6.3.
If the conjugation action of $G^{0}$ on all $\mathcal{O}(G^{m})$ is trivial then the following hold:
  1. (1)
    (36) is an isomorphism;
  2. (2)
    if $\Gamma_{1}$ is a topological group and $A$ is a topological $\mathcal{O}$-algebra, then (35) induces a bijection between continuous representations and continuous $G$-pseudocharacters;
In particular, if $G=G^{0}$ or $G^{0}$ is trivial then (36) is an isomorphism and hence $\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}}\cong\mathrm{PC}^{\Gamma_{1}}_{\underline{\Delta}}$.
Proof.
Choose $\mathcal{O}$-algebra generators $f_{1},\dots,f_{r}\in\mathcal{O}(G)$ and let $\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G}(A)$. The functions $\Theta_{1}(f_{1}),\dots,\Theta_{1}(f_{r})$ define a unique map $\rho:\Gamma_{1}\to G(A)$ such that ${\Theta_{1}(f_{1})=f_{i}\circ\rho}$ for all $i=1,\dots,r$. Let $\mu:\mathcal{O}(G)\to\mathcal{O}(G)\otimes_{\mathcal{O}}\mathcal{O}(G)$ be the comultiplication map. We can write
$\mu(f_{i})=\sum_{j,k}a_{ijk}f_{j}\otimes f_{k}$
for some $a_{ijk}\in\mathcal{O}$. By (2) of 6.1, we have
$f_{i}(\rho(\gamma_{1}\gamma_{2}))$ $=\Theta_{1}(f_{i})(\gamma_{1}\gamma_{2})=\Theta_{2}(\mu(f_{i}))(\gamma_{1},\gamma_{2})$
$=\sum_{j,k}a_{ijk}\Theta_{1}(f_{j})(\gamma_{1})\Theta_{1}(f_{k})(\gamma_{2})$
$=\sum_{j,k}a_{ijk}f_{j}(\rho(\gamma_{1}))f_{k}(\rho(\gamma_{2}))$
$=\mu(f_{i})(\rho(\gamma_{1}),\rho(\gamma_{2}))=f_{i}(\rho(\gamma_{1})\rho(\gamma_{2})),$
so $\rho$ is a homomorphism. For every $m\geq 1$, every function $f\in\mathcal{O}(G^{m})$ can be written as a linear combination of functions of the form $(g_{1},\dots,g_{m})\mapsto f_{i}(g_{j})$. So by rule (1) of 6.1 the function $\Theta_{m}(f)$ is determined by $\Theta_{1}(f_{1}),\dots,\Theta_{1}(f_{r})$. It follows, that $\Theta$ is the only $G$-pseudocharacter satisfying $\Theta_{1}(f_{i})=f_{i}\circ\rho$ for all $i$. By definition $\Theta=\Theta_{\rho}$, so this shows surjectivity and injectivity of (36).
For the claim about continuity we observe, that continuity of $\Theta_{1}(f_{1}),\dots,\Theta_{1}(f_{r})$ is equivalent to continuity of $\rho$. ∎
Corollary 6.4.
The map
(41) $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})\to\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})^{G^{0}}$
corresponding to (39) is an isomorphism.
Proof.
Via the canonical projection $\Pi:G\to\underline{\Delta}$ we have a natural $G^{0}$-action on $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})$, which is trivial, since $\underline{\Delta}^{0}=1$. So the natural $G^{0}$-equivariant map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})\to\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})$ induces a map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})\to\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})^{G^{0}}$. The representation $\pi:\Gamma_{1}\to\Delta$ corresponds to a map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})\to\mathcal{O}$. The tensor product $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$ represents $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$. By virtue of 2.3 we have a natural isomorphism
$\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})^{G^{0}}\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}\xrightarrow{\cong}\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})^{G^{0}}.$
On the other hand $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ is representable by
$\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})\otimes_{\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}\cong\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$
by 6.3. We conclude, that (41) is obtained from (40) by tensoring with $-\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$. So by 6.2 (41) is an isomorphism as well. ∎
Corollary 6.5.
Assume, there is a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. Then we have isomorphisms $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})\cong\mathcal{O}[H_{1}(\Gamma_{1},M)]\cong\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]^{G^{0}}$ via (33) and (41).
Lemma 6.6.
Assume, that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. Then the map $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})\to\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ corresponding to (38) has a section as $\mathcal{O}$-algebras. In particular, for all $A\in\mathcal{O}\text{-}\mathrm{alg}$ and all ${\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)}$, there is a representation $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$, such that $\Theta=\Theta_{\rho}$.
Proof.
Let $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$. In the diagram
[…]
the left solid square commutes, since (30) is a $G^{0}$-equivariant isomorphism. The right solid square commutes, since (33) is by 6.4 induced by $\varphi_{\mathrm{nat}}$. The section of the right vertical arrow is defined by mapping $t_{1},\dots,t_{s}$ to $1$ and since all horizontal maps are isomorphisms defines a section $\sigma$ of the left vertical arrow. Given a $G$-pseudocharacter $\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$, it corresponds to a homomorphism $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})\to A$ and by restriction along $\sigma$, we get a representation $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$, which by restriction along (41) recovers $\Theta$, so $\Theta=\Theta_{\rho}$. ∎
7. Profinite completion
So far we have worked with representations and pseudocharacters of an abstract group $\Gamma_{1}$ and topology did not play a role. In this section we transfer some of the results proved in Sections 5 and 6 to the profinite completion $\widehat{\Gamma}_{1}$ of $\Gamma_{1}$ and impose continuity conditions on the representations and pseudocharacters of $\widehat{\Gamma}_{1}$.
Let $\mathcal{O}$ be the ring of integers in a finite extension of $L$ with uniformiser $\varpi$ and residue field $k$.
7.1. Deformations of representations
Let $\overline{\rho}:\widehat{\Gamma}_{1}\rightarrow G(k)$ be a continuous representation, such that $\overline{\rho}|_{\Gamma_{1}}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(k)$.
Let $D^{\square}_{\overline{\rho}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the functor such that $D^{\square}_{\overline{\rho}}(A)$ is the set of continuous representations $\rho_{A}:\widehat{\Gamma}_{1}\rightarrow G(A)$ such that $\rho_{A}\equiv\overline{\rho}\pmod{\mathfrak{m}_{A}}$. The functor $D^{\square}_{\overline{\rho}}$ is pro-representable by $R^{\square}_{\overline{\rho}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$.
Lemma 7.1.
There is a natural isomorphism in $\widehat{\mathfrak{A}}_{\mathcal{O}}$:
[…]
where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})$ corresponding to $\overline{\rho}$ and the limit is taken over all the ideals $I$, such that the quotient is finite.
Proof.
If $A\in\mathfrak{A}_{\mathcal{O}}$ and $\rho\in D^{\square}_{\overline{\rho}}(A)$ then $\rho|_{\Gamma_{1}}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ and hence we obtain a natural homomorphism of $\mathcal{O}$-algebras $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow A$. The condition $\rho\equiv\overline{\rho}\pmod{\mathfrak{m}_{A}}$ implies that this map extends to the localisation $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})_{\mathfrak{m}}\rightarrow A$. Since $A$ is finite the kernel of this map is one of the ideals $I$ appearing in the inductive system in (42).
Conversely, if $I$ is an ideal of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})_{\mathfrak{m}}$ such that the quotient $A$ is finite then $A\in\mathfrak{A}_{\mathcal{O}}$. By specialising the universal representation $\Gamma_{1}\rightarrow G(\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}))$ along $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow A$ we obtain a representation $\rho:\Gamma_{1}\rightarrow G(A)$. Since the target is finite $\rho$ extends uniquely to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\rightarrow G(A)$, which defines a point in $D^{\square}_{\overline{\rho}}(A)$.
Hence, the right hand side of (42) pro-represents $D^{\square}_{\overline{\rho}}$ and the assertion follows.
Lemma 7.2.
Let $\mathcal{A}$ be an abelian group and let $\mathfrak{m}$ be a maximal ideal of the group ring $\mathcal{O}[\mathcal{A}]$ with residue field $k$. Then there is an isomorphism in $\widehat{\mathfrak{A}}_{\mathcal{O}}$:
[…]
where $N$ is the pro-$p$ completion of $\mathcal{A}$ and the limit is taken over all the ideals $I$, such that the quotient is finite.
Proof.
Let $\overline{\psi}:\mathcal{A}\rightarrow k^{\times}$ be the character obtained by the composition $\mathcal{A}\rightarrow\mathcal{O}[\mathcal{A}]/\mathfrak{m}=k$ and let $\psi:\mathcal{A}\rightarrow\mathcal{O}^{\times}$ be any character lifting $\overline{\psi}$ (for example the Teichmüller lift). Then $\mathfrak{m}$ is generated by $\varpi$ and $(\psi(a)^{-1}a-1)$ for all $a\in\mathcal{A}$. The map $\varphi:\mathcal{O}[\mathcal{A}]\rightarrow\mathcal{O}[\mathcal{A}]$, which sends $a\in\mathcal{A}$ to $\psi(a)a$ is an isomorphism of $\mathcal{O}$-algebras, which sends $\mathfrak{m}$ to the maximal ideal corresponding to the trivial character. We thus may assume that $\mathfrak{m}$ is generated by $\varpi$ and the augmentation ideal of $\mathcal{O}[\mathcal{A}]$.
If $P$ is a finite $p$-power order quotient of $\mathcal{A}$ then $\mathcal{O}/\varpi^{n}[P]$ is a finite local ring and the surjection $\mathcal{O}[\mathcal{A}]\twoheadrightarrow\mathcal{O}/\varpi^{n}[P]$ maps $\mathfrak{m}$ to the maximal ideal of $\mathcal{O}/\varpi^{n}[P]$. Hence, the map factors through $\mathcal{O}[\mathcal{A}]_{\mathfrak{m}}\twoheadrightarrow\mathcal{O}/\varpi^{n}[P]$. Conversely, if $A$ is a quotient of $\mathcal{O}[\mathcal{A}]_{\mathfrak{m}}$, which is finite (as a set), then the image of $\mathcal{A}$ is contained in $1+\mathfrak{m}_{A}$, and hence is a finite group of $p$-power order. This implies the assertion. ∎
Lemma 7.3.
Let $\mathcal{A}$ be an abelian group and let $\mathfrak{m}$ be a maximal ideal of the group ring $\mathcal{O}[\mathcal{A}]$ with residue field $k$ and let $N$ be the pro-$p$ completion of $\mathcal{A}$. Then the following are equivalent:
  1. (1)
    $N$ is a finitely generated $\mathbb{Z}_{p}$-module;
  2. (2)
    $\mathcal{A}/p\mathcal{A}$ is a finite $p$-group;
  3. (3)
    $\mathcal{O}\llbracket N\rrbracket$ is noetherian.
If the equivalent conditions hold then $\mathcal{O}\llbracket N\rrbracket$ is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}[\mathcal{A}]$.
Proof.
The equivalence of (1) and (2) follows from Nakayama’s lemma and the fact that $N/pN$ is the pro-$p$ completion of $\mathcal{A}/p\mathcal{A}\cong\oplus_{I}\mathbb{F}_{p}$ for some set $I$.
If $N$ is finitely generated as $\mathbb{Z}_{p}$-module, then $N\cong\mu\oplus\mathbb{Z}_{p}^{r}$, where $\mu$ is a finite $p$-group and $r=\dim_{\mathbb{Q}_{p}}N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. Hence, $\mathcal{O}\llbracket N\rrbracket\cong\mathcal{O}[\mu]\llbracket x_{1},\ldots,x_{r}\rrbracket$ and hence (1) implies (3).
If $N$ is not finitely generated as a $\mathbb{Z}_{p}$-module then we may find a strictly increasing nested family of finitely generated $\mathbb{Z}_{p}$-submodules $N_{i}\subset N$. Since they are finitely generated they are closed in $N$. The kernels of $\mathcal{O}\llbracket N\rrbracket\twoheadrightarrow\mathcal{O}\llbracket N/N_{i}\rrbracket$ form a strictly increasing nested family of ideals in $\mathcal{O}\llbracket N\rrbracket$. Hence, (3) implies (1).
As in the proof of Lemma 7.2 we may assume that $\mathfrak{m}$ is generated by $\varpi$ and the augmentation ideal of $\mathcal{O}[\mathcal{A}]$. The map $\mathcal{A}\rightarrow\mathfrak{m}$, $a\mapsto a-1$ induces an isomorphism of $k$-vector spaces $k\otimes_{\mathbb{Z}}\mathcal{A}\cong\mathfrak{m}/(\varpi,\mathfrak{m}^{2})$. If (2) holds then $\dim_{k}\mathfrak{m}/(\varpi,\mathfrak{m}^{2})$ is finite. Hence, the powers of $\mathfrak{m}$ form a cofinal system in the inverse system in (43) and the last assertion follows from Lemma 7.2. ∎
Lemma 7.4.
If $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is a finite $p$-group and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty then $R^{\square}_{\overline{\rho}}$ is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$, where $\mathfrak{m}$ is the maximal ideal corresponding to $\overline{\rho}|_{\Gamma_{1}}$.
Proof.
Proposition 5.2 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$. It follows from (17) that if $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite then $\mathcal{E}/p\mathcal{E}$ is also finite. Since $M$ is a finite free $\mathbb{Z}$-module, the assertion follows from Lemma 7.3 applied to $\mathcal{A}=(\mathcal{E}\otimes M)_{\Delta}$. ∎
Lemma 7.5.
There is a natural isomorphism of local $k$-algebras between $R^{\square}_{\overline{\rho}}/\varpi$ and the completed group algebra $k\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}.\rrbracket$.
Proof.
The ring $R^{\square}_{\overline{\rho}}/\varpi$ represents the restriction of $D^{\square}_{\overline{\rho}}$ to $\mathfrak{A}_{k}$. Lemma 7.1 implies that
[…]
Since $\overline{\rho}|_{\Gamma_{1}}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(k)$ we may apply Proposition 5.2 with the base ring $\mathcal{O}$ equal to $k$ to obtain
(45) $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})/\varpi\cong k[(\mathcal{E}\otimes M)_{\Delta}].$
The assertion then follows from (44), (45) and Lemma 7.2. ∎
Let $\Gamma\twoheadrightarrow\widehat{\Gamma}_{1}$ be a surjection of profinite groups, such that for all $A\in\mathfrak{A}_{\mathcal{O}}$ every continuous representation $\rho_{A}:\Gamma\rightarrow G(A)$ factors through $\widehat{\Gamma}_{1}$.
Let $\operatorname{ad}\overline{\rho}$ be the representation of $\Gamma$ on $\operatorname{Lie}G_{k}$ obtained by composing $\overline{\rho}$ with the adjoint representation. We write $H^{i}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$ for the $i$-th continuous group cohomology and denote by $h^{i}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$ its dimension as a $k$-vector space.
Proposition 7.6.
If $h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$ is finite then $((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$ is a finitely generated $\mathbb{Z}_{p}$-module. Moreover, if $h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$ is also finite and
(46) $h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\leq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}-\dim G_{k}$
then $R^{\square}_{\overline{\rho}}$ is complete intersection, $\mathcal{O}$-flat of relative dimension $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$.
Proof.
The assumption on $\Gamma$ implies that we could have equivalently defined the deformation problem $D^{\square}_{\overline{\rho}}$ with $\Gamma$ instead of $\widehat{\Gamma}_{1}$.
By standard obstruction theory due to Mazur, we have a presentation
(47) $\frac{\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket}{(f_{1},\ldots,f_{s})}\cong R^{\square}_{\overline{\rho}},$
where $r=\dim_{k}Z^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$ and $s=h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$. The exact sequence
$0\rightarrow(\operatorname{ad}\overline{\rho})^{\Gamma}\rightarrow\operatorname{ad}\overline{\rho}\rightarrow Z^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\rightarrow H^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\rightarrow 0$
implies that $r=h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})+\dim G_{k}$. Let $N:=((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$. By considering (47) modulo $\varpi$ and using Lemma 7.5 we deduce that $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N\leq r$. Moreover, the assumption $r-s\leq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ implies that $r-s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ and $\varpi,f_{1},\ldots,f_{s}$ can be extended to a system of parameters of a regular ring $\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket$. Thus $\varpi,f_{1},\ldots,f_{s}$ are a part of a regular sequence, and hence $R^{\square}_{\overline{\rho}}$ is complete intersection, flat over $\mathcal{O}$ of relative dimension of $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$. ∎
Corollary 7.7.
If the assumptions of 7.6 hold then there is an isomorphism of local $\mathcal{O}^{\prime}$-algebras:
$\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\square}_{\overline{\rho}}\cong\mathcal{O}^{\prime}\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}\rrbracket,$
where $\mathcal{O}^{\prime}$ is the ring of integers in a finite extension of $L$. In particular, $\operatorname{Rep}^{\widehat{\Gamma}_{1}}_{G,\pi}(\mathcal{O}^{\prime})$ and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O}^{\prime})$ are non-empty.
Proof.
It follows from Proposition 7.6 that there is a homomorphism of $\mathcal{O}$-algebras $x:R^{\square}_{\overline{\rho}}\rightarrow\overline{\mathbb{Q}}_{p}$. Then $\kappa(x)$ is a finite extension of $L$ and the image of $x$ is contained in the ring of integers of $\kappa(x)$, which we denote by $\mathcal{O}^{\prime}$. Let $\rho_{x}:\widehat{\Gamma}_{1}\rightarrow G(\mathcal{O}^{\prime})$ be the specialisation of the universal deformation along $x:R^{\square}_{\overline{\rho}}\rightarrow\mathcal{O}^{\prime}$. Then $\rho_{x}\in\operatorname{Rep}^{\widehat{\Gamma}_{1}}_{G,\pi}(\mathcal{O}^{\prime})$ and its restriction to $\Gamma_{1}$ defines a point in $\operatorname{Rep}^{\Gamma_{1}}_{G}(\mathcal{O}^{\prime})$.
Since $\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\square}_{\overline{\rho}}$ represents the functor $D^{\square}_{\overline{\rho}_{k^{\prime}}}:\mathfrak{A}_{\mathcal{O}^{\prime}}\rightarrow\operatorname{Set}$ where $\overline{\rho}_{k^{\prime}}$ is the composition of $\overline{\rho}$ with $G(k)\hookrightarrow G(k^{\prime})$, where $k^{\prime}$ is the residue field of $\mathcal{O}^{\prime}$, we may assume that $\mathcal{O}^{\prime}=\mathcal{O}$.
Proposition 5.2 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$. This identification depends on the lift $\rho_{x}$. The assertion follows from Lemmas 7.1 and 7.2. ∎
7.2. Deformations of pseudocharacters
If $A$ is a topological ring then a $G$-pseudocharacter $\Theta\in\mathrm{PC}_{G,\pi}^{\widehat{\Gamma}_{1}}(A)$ is continuous, if every $\Theta_{m}$ takes values in the subset $\mathcal{C}(\widehat{\Gamma}_{1}^{m},A)\subseteq\mathrm{Map}(\widehat{\Gamma}_{1}^{m},A)$ of continuous maps. We write $\mathrm{cPC}^{\widehat{\Gamma}_{1}}_{G,\pi}(A)\subseteq\mathrm{PC}^{\widehat{\Gamma}_{1}}_{G,\pi}(A)$ for the subset of continuous $G$-pseudocharacters.
Lemma 7.8.
Assume, that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. Let $A$ be a finite discrete $\mathcal{O}$-algebra. Then the natural map $\mathrm{cPC}^{\widehat{\Gamma}_{1}}_{G,\pi}(A)\to\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A),\Theta\mapsto\Theta|_{\Gamma_{1}}$ is bijective.
Proof.
As the image of $\Gamma_{1}$ in $\widehat{\Gamma}_{1}$ is dense, injectivity follows from [12, Lemma 3.2]. Let $\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$. By 6.6 there is a representation $\rho\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$, such that $\Theta=\Theta_{\rho}$. Since $A$ is finite, $\rho$ extends to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\to G(A)$ and we have $\Theta_{\tilde{\rho}}|_{\Gamma_{1}}=\Theta$. ∎
Let $\overline{\Theta}$ be the $G$-pseudocharacter of $\widehat{\Gamma}_{1}$ associated to $\overline{\rho}$. Let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the deformation functor which sends $A$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\widehat{\Gamma}_{1}$ valued in $A$ such that $\Theta\otimes_{A}k=\overline{\Theta}.$ The functor $D^{\mathrm{ps}}_{\overline{\Theta}}$ is pro-represented by $R^{\mathrm{ps}}_{\overline{\Theta}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ by [12, Theorem 5.4].
Proposition 7.9.
Assume, that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. Then there is an isomorphism in $\widehat{\mathfrak{A}}_{\mathcal{O}}$:
[…]
where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ corresponding to $\overline{\Theta}|_{\Gamma_{1}}$, and the limit is taken over all ideals $I$ such that the quotient is finite.
Proof.
If $A\in\mathfrak{A}_{\mathcal{O}}$ then we have natural bijections
$\operatorname{Hom}_{\widehat{\mathfrak{A}}_{\mathcal{O}}}(R^{\mathrm{ps}}_{\overline{\Theta}},A)\cong\{\Theta\in\mathrm{cPC}^{\widehat{\Gamma}_{1}}_{G,\pi}(A)\mid\Theta\otimes_{A}k=\overline{\Theta}\},$
$\operatorname{Hom}_{\text{local }\mathcal{O}\text{-}\mathrm{alg}}(\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})_{\mathfrak{m}},A)\cong\{\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)\mid\Theta\otimes_{A}k=\overline{\Theta}\}.$
As $\mathrm{cPC}^{\widehat{\Gamma}_{1}}_{G,\pi}(A)\cong\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$ by 7.8, the claim follows. ∎
Lemma 7.10.
Assume, that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. Then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is isomorphic to the completed group algebra $\mathcal{O}\llbracket H_{1}(\Gamma_{1},M)^{\wedge,p}\rrbracket$.
Proof.
Since $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty Corollary 6.5 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[H_{1}(\Gamma,M)]$. The assertion follows from Proposition 7.9 and Lemma 7.2. ∎
Lemma 7.11.
If $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is a finite $p$-group and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian and is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$, where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ corresponding to $\overline{\Theta}|_{\Gamma_{1}}$.
Proof.
The exact sequence of $\Gamma_{1}$-modules $0\rightarrow I_{\Delta}\otimes M\rightarrow\mathbb{Z}[\Delta]\otimes M\rightarrow M\rightarrow 0$ induces an exact sequence in homology:
$H_{1}(\Gamma_{2},M)\rightarrow H_{1}(\Gamma_{1},M)\rightarrow(I_{\Delta}\otimes M)_{\Delta}.$
Since the action of $\Gamma_{2}$ on $M$ is trivial we have a canonical isomorphism $H_{1}(\Gamma_{2},M)\cong\Gamma_{2}^{\mathrm{ab}}\otimes M$. Since $I_{\Delta}$ and $M$ are free $\mathbb{Z}$-modules of finite rank we conclude that the assumption $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite implies that $\mathcal{A}/p\mathcal{A}$ is finite, where $\mathcal{A}=H_{1}(\Gamma_{1},M)$. The assertion follows from Lemmas 7.10 and 7.3. ∎
7.3. Moduli space of representations
We study the following schemes.
Definition 7.12.
For $\mathcal{G}=\Gamma_{1}$ or $\widehat{\Gamma}_{1}$, let $X^{\mathrm{gen},\mathcal{G}}_{G,\overline{\Theta}}:R^{\mathrm{ps}}_{\overline{\Theta}}\text{-}\mathrm{alg}\rightarrow\operatorname{Set}$ be the functor
$X^{\mathrm{gen},\mathcal{G}}_{G,\overline{\Theta}}(A):=\{\rho\in\operatorname{Rep}^{\mathcal{G}}_{G,\pi}(A):\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A\},$
where $\Theta^{u}\in D^{\mathrm{ps}}_{\overline{\Theta}}(R^{\mathrm{ps}}_{\overline{\Theta}})$ is the universal deformation of $\overline{\Theta}$ and we consider its restriction to $\Gamma_{1}$ if $\mathcal{G}=\Gamma_{1}$.
Proposition 7.13.
Assume that $\operatorname{Rep}^{\widehat{\Gamma}_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. The restriction to $\Gamma_{1}$ induces an isomorphism
$X^{\mathrm{gen},\widehat{\Gamma}_{1}}_{G,\overline{\Theta}}\overset{\cong}{\longrightarrow}X^{\mathrm{gen},\Gamma_{1}}_{G,\overline{\Theta}}.$
In particular, $X^{\mathrm{gen},\widehat{\Gamma}_{1}}_{G,\overline{\Theta}}$ is representable by the $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra isomorphic to
$R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})}\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
where $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
Proof.
We let $\mathcal{G}$ be either $\Gamma_{1}$ or its profinite completion $\widehat{\Gamma}_{1}$ in the proof. The functor $X^{\mathrm{gen},\mathcal{G}}_{G,\overline{\Theta}}$ is representable by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}(\mathrm{PC}^{\mathcal{G}}_{G,\pi})}\mathcal{O}(\operatorname{Rep}^{\mathcal{G}}_{G,\pi})$. Let $\rho_{0}\in\operatorname{Rep}^{\widehat{\Gamma}_{1}}_{G,\pi}(\mathcal{O})$ then $\rho_{0}|_{\Gamma_{1}}$ is in $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$, and using these representations we may identify $\operatorname{Rep}^{\mathcal{G}}_{G,\pi}$ with the space of $1$-cocycles $Z^{1}(\mathcal{G},\mathfrak{D}(M)(-))$ by Proposition 5.2. Corollaries 4.11 and 6.5 imply that after choosing a basis of $I_{\Delta}M$ as a $\mathbb{Z}$-module we may identify
$\mathcal{O}(\operatorname{Rep}^{\mathcal{G}}_{G,\pi})\cong\mathcal{O}(\mathrm{PC}^{\mathcal{G}}_{G,\pi})[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
thus for both $\mathcal{G}=\Gamma_{1}$ and $\mathcal{G}=\widehat{\Gamma}_{1}$ we have
(49) $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}(\mathrm{PC}^{\mathcal{G}}_{G,\pi})}\mathcal{O}(\operatorname{Rep}^{\mathcal{G}}_{G,\pi})\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$
and under theses isomorphisms the restriction to $\Gamma_{1}$ is just the identity map. ∎
The following Lemma will allow us to relate the scheme $X^{\mathrm{gen},\widehat{\Gamma}_{1}}_{G,\overline{\Theta}}$ to the scheme $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ introduced in [11].
Lemma 7.14.
Let $\tau:G\hookrightarrow\mathbb{A}^{n}$ be a closed immersion of $\mathcal{O}$-schemes, let $A$ be an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra and let $\rho\in X^{\mathrm{gen},\widehat{\Gamma}_{1}}_{G,\overline{\Theta}}(A)$. Assume that $\operatorname{Rep}^{\widehat{\Gamma}_{1}}_{G,\pi}(\mathcal{O})$ is non-empty. Then $\tau(\rho(\widehat{\Gamma}_{1}))$ is contained in a finitely generated $R^{\mathrm{ps}}_{\overline{\Theta}}$-submodule of $A^{n}=\mathbb{A}^{n}(A)$.
Proof.
Since $X^{\mathrm{gen},\widehat{\Gamma}_{1}}_{G,\overline{\Theta}}$ is represented by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}(\mathrm{PC}^{\widehat{\Gamma}_{1}}_{G,\pi})}\mathcal{O}(\operatorname{Rep}^{\widehat{\Gamma}_{1}}_{G,\pi})$ the assertion follows from Lemma 5.7 applied to $\widehat{\Gamma}_{1}$, and 6.5. ∎
7.4. Irreducible components
Let $N$ be a finitely generated $\mathbb{Z}_{p}$-module and let $\mathcal{O}\llbracket N\rrbracket$ be the completed group algebra of $N$, where $\mathcal{O}$ is the ring of integers in a finite extension $L$ of $\mathbb{Q}_{p}$. Let $\widehat{\mathfrak{D}}(N):\widehat{\mathfrak{A}}_{\mathcal{O}}\rightarrow\operatorname{Ab}$ be a formal group scheme, given by
[…]
Multiplication in $\widehat{\mathfrak{D}}(N)$ is induced by the map
(51) $c:\mathcal{O}\llbracket N\rrbracket\rightarrow\mathcal{O}\llbracket N\rrbracket\operatorname{\widehat{\otimes}}_{\mathcal{O}}\mathcal{O}\llbracket N\rrbracket,\quad n\mapsto n\operatorname{\widehat{\otimes}}n,\quad\forall n\in N.$
Let $\mu$ be the torsion subgroup of $N$. Then we have a non-canonical isomorphism $N\cong\mu\oplus\mathbb{Z}_{p}^{r}$, where $r=\dim_{\mathbb{Q}_{p}}N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$. This induces an isomorphism
(52) $\mathcal{O}\llbracket N\rrbracket\cong\mathcal{O}[\mu]\llbracket x_{1},\ldots,x_{r}\rrbracket.$
We assume that $L$ contains all the $p^{m}$-th roots of unity, where $p^{m}$ is the order of $\mu$. Then the group $\mathrm{X}(\mu)$ of characters $\chi:\mu\rightarrow\mathcal{O}^{\times}$ also has order $p^{m}$. The following Lemma is an immediate consequence of (52):
Lemma 7.15.
The following hold:
  1. (1)
    the irreducible components of $\operatorname{Spec}\mathcal{O}\llbracket N\rrbracket$ are in canonical bijection with $\mathrm{X}(\mu)$, so that the irreducible component corresponding to $\chi\in\mathrm{X}(\mu)$ is given by $\operatorname{Spec}(\mathcal{O}\llbracket N\rrbracket\otimes_{\mathcal{O}[\mu],\chi}\mathcal{O})$;
  2. (2)
    every irreducible component of $\operatorname{Spec}\mathcal{O}\llbracket N\rrbracket$ contains an $\mathcal{O}$-rational point $\psi\in\widehat{\mathfrak{D}}(N)(\mathcal{O})$;
  3. (3)
    a point $\psi\in\widehat{\mathfrak{D}}(N)(\mathcal{O})=\operatorname{Hom}^{\mathrm{cont}}_{\operatorname{Group}}(N,\mathcal{O}^{\times})$ lies on the irreducible component corresponding to $\chi\in\mathrm{X}(\mu)$ if and only if $\psi(x)=\chi(x)$ for all $x\in\mu$.
Let $R$ be a complete local noetherian $\mathcal{O}$-algebra with residue field $k$. Let ${\mathcal{X}:\widehat{\mathfrak{A}}_{\mathcal{O}}\rightarrow\operatorname{Set}}$ be the functor $\mathcal{X}(A)=\operatorname{Hom}_{\widehat{\mathfrak{A}}_{\mathcal{O}}}(R,A)$. This functor is represented by a formal scheme $\operatorname{Spf}R$. Let us suppose that we have a faithful and transitive action of $\widehat{\mathfrak{D}}(N)$ on $\mathcal{X}$. Concretely, this means that for all $A\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ we have a faithful and transitive action of the group $\widehat{\mathfrak{D}}(N)(A)$ on the set $\mathcal{X}(A)$, which is functorial in $A$. It is enough to restrict to $A\in\mathfrak{A}_{\mathcal{O}}$ as the general case follows by continuity. The action map $\widehat{\mathfrak{D}}(N)\times\mathcal{X}\rightarrow\mathcal{X}$ gives us a morphism in $\widehat{\mathfrak{A}}_{\mathcal{O}}$:
(53) $\alpha:R\rightarrow\mathcal{O}\llbracket N\rrbracket\operatorname{\widehat{\otimes}}_{\mathcal{O}}R.$
Let us assume that $\mathcal{X}(\mathcal{O})$ is non-empty and choose $x\in\mathcal{X}(\mathcal{O})$. Since the action map is faithful and transitive for every $A\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ the map
(54) $\widehat{\mathfrak{D}}(N)(A)\rightarrow\mathcal{X}(A),\quad\psi\mapsto\psi\cdot x_{A}$
is bijective, where $x_{A}$ is the image of $x$ in $\mathcal{X}(A)$. This implies that the composition
[…]
is an isomorphism.
Example 7.16.
If $\mathcal{X}=\widehat{\mathfrak{D}}(N)$ and the action is given by left translations, then it follows from (50), (51) that $\alpha_{\psi}$ corresponds to the character $N\rightarrow\mathcal{O}\llbracket N\rrbracket^{\times}$, $n\mapsto\psi(n)n$.
Lemma 7.17.
Assume that $\mathcal{X}(\mathcal{O})$ is non-empty. Then the following hold:
  1. (1)
    every irreducible component of $\operatorname{Spec}R$ has an $\mathcal{O}$-rational point $x\in\mathcal{X}(\mathcal{O})$;
  2. (2)
    every $x\in\mathcal{X}(\mathcal{O})$ lies on a unique irreducible component of $\operatorname{Spec}R$;
  3. (3)
    if $x,y\in\mathcal{X}(\mathcal{O})$ then there exists a unique $\psi\in\widehat{\mathfrak{D}}(N)(\mathcal{O})$ such that $\psi\cdot x=y$;
  4. (4)
    let $x$, $y$ and $\psi$ be as in part $(3)$. Then $x$, $y$ lie on the same irreducible component of $\operatorname{Spec}R$ if and only if $\psi|_{\mu}=1$.
Proof.
If $\mathcal{X}=\widehat{\mathfrak{D}}(N)$ with the action given by left translations then the assertions follow readily from 7.15.
In the general case, we pick $x\in\mathcal{X}(\mathcal{O})$. The isomorphism $\alpha_{x}:R\rightarrow\mathcal{O}\llbracket N\rrbracket$ of $\mathcal{O}$-algebras allows us to reduce the question to the case above. For parts (3) and (4), we note that the bijection (54) is $\widehat{\mathfrak{D}}(N)(A)$-equivariant for the action by left translations on the source. ∎
Lemma 7.18.
Assume that $\mathcal{X}(\mathcal{O})$ is non-empty. Then the action of $\widehat{\mathfrak{D}}(N)(\mathcal{O})$ on $\mathcal{X}(\mathcal{O})$ induces a canonical action of $\mathrm{X}(\mu)$ on the set of irreducible components of $\operatorname{Spec}R$. This action is faithful and transitive.
Proof.
Since $\mu$ is a direct summand of $N$ the map $\psi\mapsto\psi|_{\mu}$ induces a surjective group homomorphism $\widehat{\mathfrak{D}}(N)(\mathcal{O})\twoheadrightarrow\mathrm{X}(\mu)$. Let $K$ be the kernel of this map. 7.17 implies that there is a natural bijection between the set of irreducible components of $\operatorname{Spec}R$ and the set of $K$-orbits in $\mathcal{X}(\mathcal{O})$. The action $\widehat{\mathfrak{D}}(N)(\mathcal{O})$ on the set of $K$-orbits factors through the action of $\mathrm{X}(\mu)$, which induces the sought after action on the set of irreducible components. Since the action of $\widehat{\mathfrak{D}}(N)(\mathcal{O})$ on $\mathcal{X}(\mathcal{O})$ is faithful and transitive the same applies to the action of $\mathrm{X}(\mu)$. ∎
We will now get back to our particular example. Let $\widehat{Z}^{1}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Ab}$ be the functor such that $\widehat{Z}^{1}(A)$ is the set of continuous $1$-cocycles $\Phi:\widehat{\Gamma}_{1}\rightarrow\operatorname{Hom}(M,1+\mathfrak{m}_{A})$ for the discrete topology on the target.
Lemma 7.19.
The functor $\widehat{Z}^{1}$ is pro-represented by $\mathcal{O}\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}\rrbracket$.
Proof.
This is a special case of Lemma 7.2 applied with $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k),\gamma\mapsto(1,\pi(\gamma))$ via Proposition 5.2 with $\rho_{0}:\Gamma_{F}\rightarrow G(\mathcal{O}),\gamma\mapsto(1,\pi(\gamma))$. ∎
It follows from Proposition 5.2 that if $\Phi\in\widehat{Z}^{1}(A)$ and $\rho\in D^{\square}_{\overline{\rho}}(A)$ then $\gamma\mapsto\Phi(\gamma)\rho(\gamma)$ defines a representation $\Phi\rho\in D^{\square}_{\overline{\rho}}(A)$ and the map $\widehat{Z}^{1}\times D^{\square}_{\overline{\rho}}\rightarrow D^{\square}_{\overline{\rho}}$, $(\Phi,\rho)\mapsto\Phi\rho$ defines a faithful and transitive action of $\widehat{Z}^{1}$ on $D^{\square}_{\overline{\rho}}$.
Proposition 7.20.
Assume that $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite and let $\mu$ be the torsion subgroup of $((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$. Assume further that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty and $\mathcal{O}$ contains all the $p^{m}$-th roots of unity, where $p^{m}$ is the order of $\mu$. Then there is a canonical action of the character group $\mathrm{X}(\mu)$ on the set of irreducible components of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$, $\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ and $X^{\mathrm{gen}}_{G,\overline{\Theta}}$, respectively. Moreover, the following hold:
  1. (1)
    this action is faithful and transitive;
  2. (2)
    irreducible components of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ and $\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ are formally smooth over $\mathcal{O}$;
  3. (3)
    irreducible components of $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ are of the form
    $\operatorname{Spec}\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
    where $r=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}H_{1}(\Gamma_{1},M)^{\wedge,p}$ and $r+s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$.
Proof.
The assumption that $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite implies that $((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$ is a finitely generated $\mathbb{Z}_{p}$-module by 7.3, and hence $\mu$ is a finite $p$-group.
The assumption that $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty implies via Proposition 5.2 that $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\cong\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$. It follows from Lemmas 7.1,, and 7.2 that $R^{\square}_{\overline{\rho}}\cong\mathcal{O}\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}\rrbracket$. In particular, $D^{\square}_{\overline{\rho}}(\mathcal{O})$ is non-empty. It follows from Lemma 7.18 that the action of $\widehat{Z}^{1}(\mathcal{O})$ on $D^{\square}_{\overline{\rho}}(\mathcal{O})$ induces a transitive and faithful action of $\mathrm{X}(\mu)$ on the set of irreducible components of $R^{\square}_{\overline{\rho}}$.
Let us spell out Lemma 7.17 in our context. Given an irreducible component $X$ of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$, $X(\mathcal{O})$ is non-empty and we pick any $\rho\in X(\mathcal{O})$; given $\chi\in\mathrm{X}(\mu)$ we pick any $\Phi\in\widehat{Z}^{1}(\mathcal{O})$ such that $\Phi(\gamma)=\chi(\gamma)$ for all $\gamma\in\mu$. Then $\chi\cdot X$ is the unique irreducible component of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ such that $\Phi\rho\in(\chi\cdot X)(\mathcal{O})$.
It follows from 4.9 that
$((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}\cong H_{1}(\Gamma_{1},M)^{\wedge,p}\oplus\mathbb{Z}_{p}^{s}$
for some $s\geq 0$. Hence, $R^{\square}_{\overline{\rho}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}\llbracket t_{1},\ldots,t_{s}\rrbracket$ and $\mu$ is the torsion subgroup of $H_{1}(\Gamma_{1},M)^{\wedge,p}$. Thus the map $\operatorname{Spec}R^{\square}_{\overline{\rho}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces an $\mathrm{X}(\mu)$-equivariant bijection between the irreducible components.
Since $A^{\mathrm{gen}}_{G,\overline{\Theta}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$ by 7.13 the map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces a an $\mathrm{X}(\mu)$-equivariant bijection between the sets of irreducible components.
The isomorphism $R^{\mathrm{ps}}_{\overline{\Theta}}\cong\mathcal{O}\llbracket H_{1}(\Gamma,M)^{\wedge,p}\rrbracket$ allows us to consider $R^{\mathrm{ps}}_{\overline{\Theta}}$ as an $\mathcal{O}[\mu]$-algebra. This is non-canonical: it amounts to distinguishing one irreducible component of $\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$. Once we do this the other irreducible are given by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}[\mu],\chi}\mathcal{O}$ for $\chi\in\mathrm{X}(\mu)$, and the special component corresponds to the trivial character. These are isomorphic to $\mathcal{O}\llbracket H_{1}(\Gamma,M)^{\wedge,p}/\mu\rrbracket$ and hence are formally smooth.
Similarly irreducible components of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ and $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ are given by
$R^{\square}_{\overline{\rho}}\otimes_{\mathcal{O}[\mu],\chi}\mathcal{O}\cong\mathcal{O}\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}/\mu\rrbracket$
and
$A^{\mathrm{gen}}_{G,\overline{\Theta}}\otimes_{\mathcal{O}[\mu],\chi}\mathcal{O}\cong\mathcal{O}\llbracket H_{1}(\Gamma,M)^{\wedge,p}/\mu\rrbracket[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
respectively. ∎
8. Rank calculations
[…]
If $\mathcal{A}$ is an abelian group we denote its pro-$p$ completion by $\mathcal{A}^{\wedge,p}$. If $N$ is a $\mathbb{Z}_{p}$-module we define $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N:=\dim_{\mathbb{Q}_{p}}(N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$.
Let $\Gamma_{E}^{\mathrm{ab}}$ be the maximal abelian pro-finite quotient of $\Gamma_{E}$ and let $\Gamma_{E}^{\mathrm{ab},p}$ be the maximal abelian pro-$p$ quotient of $\Gamma_{E}$. The Artin map $\operatorname{Art}_{E}:E^{\times}\rightarrow\Gamma_{E}^{\mathrm{ab}}$ induces an isomorphism between the profinite completion of $E^{\times}$ and $\Gamma_{E}^{\mathrm{ab}}$. Thus $(E^{\times}\otimes M)^{\wedge,p}\cong\Gamma_{E}^{\mathrm{ab},p}\otimes M$.
Lemma 8.1.
Let $N$ be a finitely generated $\mathbb{Z}[\Delta]$ (resp. $\mathbb{Z}_{p}[\Delta]$) module. Then the Tate cohomology groups $\widehat{H}^{i}(\Delta,N)$ are finite for all $i\in\mathbb{Z}$.
Proof.
If $N$ is finitely generated over $\mathbb{Z}$ then the statement is proved in [4, Corollary 2, p. 105]. The same argument carries over if $N$ is finitely generated over $\mathbb{Z}_{p}$: the cohomology groups are finitely generated $\mathbb{Z}_{p}$-modules, since the complex computing cohomology consists of finitely generated $\mathbb{Z}_{p}$-modules, moreover they are annihilated by the order of $\Delta$. ∎
Lemma 8.2.
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(M^{\wedge,p}\otimes I_{\Delta})_{\Delta}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}.$
Proof.
The long exact sequence in homology attached to
$0\rightarrow M^{\wedge,p}\otimes I_{\Delta}\rightarrow M^{\wedge,p}\otimes\mathbb{Z}[\Delta]\rightarrow M^{\wedge,p}\rightarrow 0$
yields an exact sequence
$H_{1}(\Delta,M^{\wedge,p})\rightarrow(M^{\wedge,p}\otimes I_{\Delta})_{\Delta}\rightarrow M^{\wedge,p}\rightarrow(M^{\wedge,p})_{\Delta}\rightarrow 0,$
as $M\otimes\mathbb{Z}[\Delta]$ is induced by the projection formula. Since we may swap coinvariants with completions and $M_{\Delta}$ is finitely generated we have
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(M^{\wedge,p})_{\Delta}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(M_{\Delta})^{\wedge,p}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}.$
Since $M^{\wedge,p}$ is finitely generated over $\mathbb{Z}_{p}$ and $\Delta$ is finite, the group is $H_{1}(\Delta,M^{\wedge,p})$ is finite by Lemma 8.1. This implies the assertion. ∎
Lemma 8.3.
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((E^{\times}\otimes M)^{\Delta})^{\wedge,p}.$
Proof.
Since $M$ is a free $\mathbb{Z}$-module, we have an exact sequence of $\Delta$-modules
$0\rightarrow(1+\mathfrak{p}_{E})\otimes M\rightarrow E^{\times}\otimes M\rightarrow(E^{\times}/(1+\mathfrak{p}_{E}))\otimes M\rightarrow 0$
Since $1+\mathfrak{p}_{E}$ is a finitely generated $\mathbb{Z}_{p}$-module and $E^{\times}/(1+\mathfrak{p}_{E})$ is a finitely generated $\mathbb{Z}$-module, we deduce that $\widehat{H}^{i}(\Delta,E^{\times}\otimes M)$ are finite for all $i\in\mathbb{Z}$. From the exact sequence:
$0\rightarrow\widehat{H}^{-1}(\Delta,E^{\times}\otimes M)\rightarrow(E^{\times}\otimes M)_{\Delta}\rightarrow(E^{\times}\otimes M)^{\Delta}\rightarrow\widehat{H}^{0}(\Delta,E^{\times}\otimes M)\rightarrow 0$
we deduce that the pro-$p$ completions of $(E^{\times}\otimes M)_{\Delta}$ and of $(E^{\times}\otimes M)^{\Delta}$ have the same $\mathbb{Z}_{p}$-rank. It follows from the universal property of pro-$p$ completion, that it commutes with taking $\Delta$-coinvariants. Hence the pro-$p$ completion of $(E^{\times}\otimes M)_{\Delta}$ is isomorphic to $((E^{\times})^{\wedge,p}\otimes M)_{\Delta}\cong(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}$. ∎
Lemma 8.4.
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot[F:\mathbb{Q}_{p}]+\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
Proof.
We consider the long exact sequence in homology
$H_{1}(\Delta,M^{\wedge,p})\to((\mathcal{O}_{E}^{\times})^{\wedge,p}\otimes M)_{\Delta}\to((E^{\times})^{\wedge,p}\otimes M)_{\Delta}\to(M^{\wedge,p})_{\Delta}\to 0.$
Since completion commutes with taking $\Delta$-coinvariants we have
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(M^{\wedge,p})_{\Delta}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}.$
Since $H_{1}(\Delta,M^{\wedge,p})$ is finite by 8.1, we are left to compute the $\mathbb{Z}_{p}$-rank of $((\mathcal{O}_{E}^{\times})^{\wedge,p}\otimes M)_{\Delta}$. We note that $(\mathcal{O}_{E}^{\times})^{\wedge,p}$ is equal to $1+\mathfrak{p}_{E}$ and another application of 8.1 shows that the rank does not change if we replace $(\mathcal{O}_{E}^{\times})^{\wedge,p}$ with any open $\Delta$-stable subgroup $V$ of $1+\mathfrak{p}_{E}$.
We choose $V$ to be the image of a $p$-adic exponential function defined on $\mathfrak{p}_{E}^{n}$ for some large enough $n\geq 1$. We then have an isomorphism $\mathfrak{p}_{E}^{n}\otimes M\cong V\otimes M$ of $\Delta$-modules. Since $\mathfrak{p}_{E}^{n}\cong\mathcal{O}_{E}$ is isomorphic to $\mathcal{O}_{F}[\Delta]$ as $\mathbb{Z}_{p}[\Delta]$-module, see the proof of [13, Section 1.4], $\mathcal{O}_{E}\otimes M$ is free and thus $(\mathcal{O}_{E}\otimes M)_{\Delta}\cong\mathcal{O}_{F}\otimes M$. Thus $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{O}_{E}^{\times})^{\wedge,p}\otimes M)_{\Delta}=[F:\mathbb{Q}_{p}]\cdot\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M$ and the assertion follows. ∎
Lemma 8.5.
The torsion subgroup of $((E^{\times}\otimes M)^{\Delta})^{\wedge,p}$ is equal to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$.
Proof.
The image of $(E^{\times}\otimes M)^{\Delta}\rightarrow(E^{\times}/\mathcal{O}_{E}^{\times})\otimes M$ is a free $\mathbb{Z}$-module of finite rank, which we denote by $s$, as the target is a free $\mathbb{Z}$-module of finite rank. The kernel of this map is equal to $(\mathcal{O}_{E}^{\times}\otimes M)^{\Delta}$. The Teichmüller lift gives an isomorphism of $\Delta$-modules $\mathcal{O}_{E}^{\times}\cong(1+\mathfrak{p}_{E})\oplus k_{E}^{\times}$ and hence we have an isomorphism of abelian groups
(55) $(E^{\times}\otimes M)^{\Delta}\cong((1+\mathfrak{p}_{E})\otimes M)^{\Delta}\oplus(k_{E}^{\times}\otimes M)^{\Delta}\oplus\mathbb{Z}^{s}.$
Since the order of $k_{E}^{\times}$ is prime to $p$ and $((1+\mathfrak{p}_{E})\otimes M)^{\Delta}$ is closed in $(1+\mathfrak{p}_{E})\otimes M$ and hence $p$-adically complete, we conclude that the torsion subgroup in the pro-$p$ completion of $(E^{\times}\otimes M)^{\Delta}$ coincides with the torsion subgroup in $((1+\mathfrak{p}_{E})\otimes M)^{\Delta}$, which is equal to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$. ∎
Lemma 8.6.
If $\mathcal{A}$ is a finitely generated $\mathbb{Z}[\Delta]$-module then $(\mathcal{A}^{\wedge,p})^{\Delta}\cong(\mathcal{A}^{\Delta})^{\wedge,p}$.
Proof.
We have $\mathcal{A}^{\wedge,p}\cong\mathcal{A}\otimes\mathbb{Z}_{p}$ and $(\mathcal{A}^{\Delta})^{\wedge,p}\cong\mathcal{A}^{\Delta}\otimes\mathbb{Z}_{p}$ by [14, Tag 00MA]. We may express $\mathcal{A}^{\Delta}$ as the kernel of
$\mathcal{A}\rightarrow\bigoplus_{\delta\in\Delta}\mathcal{A},\quad a\mapsto(\delta a-a)_{\delta\in\Delta}.$
Since $\mathbb{Z}_{p}$ is a flat $\mathbb{Z}$-module we conclude that $\mathcal{A}^{\Delta}\otimes\mathbb{Z}_{p}\cong(\mathcal{A}\otimes\mathbb{Z}_{p})^{\Delta}.$
Lemma 8.7.
$((E^{\times}\otimes M)^{\Delta})^{\wedge,p}\cong(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}$.
Proof.
Let $n_{0}$ be an integer such that $\exp:\mathfrak{p}_{E}^{n}\rightarrow 1+\mathfrak{p}_{E}^{n}$ converges for all $n\geq n_{0}$ and let $V_{n}:=\exp(\mathfrak{p}_{E}^{n})$. Then $V_{n}$ for $n\geq n_{0}$ form a basis of open neighbourhoods of $1$ in $1+\mathfrak{p}_{E}$. Since $\mathfrak{p}_{E}^{n}\cong\mathcal{O}_{E}\cong\mathcal{O}_{F}[\Delta]$ as $\Delta$-module and $\exp$ is $\Delta$-equivariant we have an isomorphism $V_{n}\otimes M\cong\operatorname{Ind}^{\Delta}_{1}(\mathcal{O}_{F}\otimes M)$ and hence $H^{1}(\Delta,V_{n}\otimes M)=0$. Thus for all $n\geq n_{0}$ we obtain an exact sequence
(56) $0\rightarrow(V_{n}\otimes M)^{\Delta}\rightarrow(E^{\times}\otimes M)^{\Delta}\rightarrow((E^{\times}/V_{n})\otimes M)^{\Delta}\rightarrow 0$
[…]
The isomorphism $(E^{\times}\otimes M)^{\wedge,p}\cong\Gamma_{E}^{\mathrm{ab},p}\otimes M$ induces an isomorphism
(58) $((E^{\times}/V_{n})\otimes M)^{\wedge,p}\cong(\Gamma_{E}^{\mathrm{ab},p}/\operatorname{Art}_{E}(V_{n}))\otimes M.$
Since $H^{1}(\Delta,V_{n}\otimes M)=0$ we have an exact sequence
(59) $0\rightarrow(V_{n}\otimes M)^{\Delta}\rightarrow(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}\rightarrow((\Gamma_{E}^{\mathrm{ab},p}/\operatorname{Art}_{E}(V_{n}))\otimes M)^{\Delta}\rightarrow 0$
We thus have
[…]
and the assertion follows from (58) and (57). ∎
Corollary 8.8.
There is an isomorphism of $\mathbb{Z}_{p}$-modules:
$(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}\cong(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}\times\mathbb{Z}_{p}^{r},$
where $r=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot[F:\mathbb{Q}_{p}]+\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
Proof.
This follows from Lemmas 8.7, 8.5, 8.4, 8.3. ∎
Proposition 8.9.
Let $0\rightarrow E^{\times}\rightarrow\mathcal{E}\rightarrow I_{\Delta}\rightarrow 0$ be any extension of $\mathbb{Z}[\Delta]$-modules.
(60) $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}=([F:\mathbb{Q}_{p}]+1)\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M.$
Proof.
Since $I_{\Delta}\otimes M$ is a free $\mathbb{Z}$-module the surjection $\mathcal{E}\otimes M\twoheadrightarrow I_{\Delta}\otimes M$ has a section, and hence we have an exact sequence of $\mathbb{Z}_{p}[\Delta]$-modules
$0\rightarrow(E^{\times}\otimes M)^{\wedge,p}\rightarrow(\mathcal{E}\otimes M)^{\wedge,p}\rightarrow(I_{\Delta}\otimes M)^{\wedge,p}\rightarrow 0$
Since $M$ is a free $\mathbb{Z}$-module of finite rank and $(E^{\times})^{\wedge,p}\cong\Gamma^{\mathrm{ab},p}_{E}$ we $(E^{\times}\otimes M)^{\wedge,p}\cong(E^{\times})^{\wedge,p}\otimes M\cong\Gamma^{\mathrm{ab},p}_{E}\otimes M$. Since $I_{\Delta}$ is a free $\mathbb{Z}$-module $(I_{\Delta}\otimes M)^{\wedge,p}\cong I_{\Delta}\otimes M^{\wedge,p}$, so that we obtain an exact sequence of $\mathbb{Z}_{p}[\Delta]$-modules:
$0\rightarrow\Gamma_{E}^{\mathrm{ab},p}\otimes M\rightarrow(\mathcal{E}\otimes M)^{\wedge,p}\rightarrow I_{\Delta}\otimes M^{\wedge,p}\rightarrow 0$
Taking $\Delta$-coinvariants and observing that $H_{1}(\Delta,I_{\Delta}\otimes M^{\wedge,p})\cong H_{2}(\Delta,M^{\wedge,p})$ we obtain an exact sequence:
(61) $H_{2}(\Delta,M^{\wedge,p})\rightarrow(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}\rightarrow((\mathcal{E}\otimes M)^{\wedge,p})_{\Delta}\rightarrow(I_{\Delta}\otimes M^{\wedge,p})_{\Delta}\rightarrow 0.$
Since $H_{2}(\Delta,M^{\wedge,p})$ is a torsion module by Lemma 8.1, we deduce that
$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)^{\wedge,p})_{\Delta}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(\Gamma_{E}^{\mathrm{ab},p}\otimes M)_{\Delta}+\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(I_{\Delta}\otimes M^{\wedge,p})_{\Delta}.$
The assertion (60) follows from Lemmas 8.2 and 8.4. ∎
9. Galois deformations
We will apply the machinery developed in the previous sections in an arithmetic situation. Let $F$ be a finite extension of $\mathbb{Q}_{p}$. We fix an algebraic closure $\overline{F}$ and let $\Gamma_{F}:=\operatorname{Gal}(\overline{F}/F)$. Let $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ be a continuous representation, where $G$ is a generalised torus over $\mathcal{O}$ and let $\overline{\Theta}$ be the $G$-pseudocharacter associated to $\overline{\rho}$. Let $\Gamma_{E}$ be the kernel of the composition $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\rightarrow(G/G^{0})(k)$ and let $\Delta:=\operatorname{Gal}(E/F)$. Let $\pi:\Gamma_{F}\rightarrow\Delta$ be the projection map.
We let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the functor
$D^{\mathrm{ps}}_{\overline{\Theta}}(A)=\{\Theta\in\mathrm{cPC}^{\Gamma_{F}}_{G,\pi}(A):\Theta\otimes_{A}k=\overline{\Theta}\}$
and let $R^{\mathrm{ps}}_{\overline{\Theta}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ be the ring pro-representing $D^{\mathrm{ps}}_{\overline{\Theta}}$ and let $\Theta^{u}$ be the universal deformation of $\overline{\Theta}$.
We let $X^{\mathrm{gen}}_{G,\overline{\Theta}}:R^{\mathrm{ps}}_{\overline{\Theta}}\text{-}\mathrm{alg}\rightarrow\operatorname{Set}$ be the functor
$X^{\mathrm{gen}}_{G,\overline{\Theta}}(A)=\{\rho\in\operatorname{Rep}^{\Gamma_{F}}_{G,\pi}(A):\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A\}.$
Let $A^{\mathrm{gen}}_{G,\overline{\Theta}}$ be the $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra representing $X^{\mathrm{gen}}_{G,\overline{\Theta}}$.
We let $D^{\square}_{\overline{\rho}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the functor such that $D^{\square}_{\overline{\rho}}(A)$ is the set of continuous representations $\rho:\Gamma_{F}\rightarrow G(A)$ such that $\rho\equiv\overline{\rho}\mod{\mathfrak{m}_{A}}$. Let $R^{\square}_{\overline{\rho}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ be the the ring pro-representing $D^{\square}_{\overline{\rho}}$.
The Weil group $W_{E/F}$ fits into a short exact sequence
$0\rightarrow E^{\times}\rightarrow W_{E/F}\rightarrow\Delta\rightarrow 0,$
corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$ by [15, (1.2)]. Let $\widehat{W}_{E/F}$ be the profinite completion of $W_{E/F}$. The Artin map $\operatorname{Art}_{E}:E^{\times}\rightarrow\Gamma_{E}^{\mathrm{ab}}$ induces an isomorphism between the profinite completion of $E^{\times}$ and $\Gamma_{E}^{\mathrm{ab}}$. This yields a natural isomorphism
(62) $\widehat{W}_{E/F}\cong\Gamma_{F}/[\Gamma_{E},\Gamma_{E}],$
where $[\Gamma_{E},\Gamma_{E}]$ is the closure of the subgroup generated by the commutators in $\Gamma_{E}$.
Lemma 9.1.
Every continuous representation $\rho:\Gamma_{F}\rightarrow G(A)$, where $A$ is a finite discrete $\mathcal{O}$-algebra, factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$.
Proof.
Since $\rho(\Gamma_{E})\subseteq G^{0}(A)$ and $G^{0}(A)$ is commutative and Hausdorff the assertion follows from (62). ∎
Lemma 9.2.
Let $A\in R^{\mathrm{ps}}_{\overline{\Theta}}\text{-}\mathrm{alg}$. Then every $\rho\in X^{\mathrm{gen}}_{G,\overline{\Theta}}(A)$ factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$.
Proof.
The restriction of $\overline{\rho}$ to $\Gamma_{E}$ takes values in $G^{0}(k)$. Let $\overline{\Psi}$ be the $G^{0}$-pseudocharacter of $\overline{\rho}|_{\Gamma_{E}}$ and let $\Psi^{u}\in D^{\mathrm{ps}}_{\overline{\Psi}}(R^{\mathrm{ps}}_{\overline{\Psi}})$ be the universal deformation of $\overline{\Psi}$. Since $G^{0}$ is commutative, Lemma 6.3 implies that there is a continuous group homomorphism $\psi^{u}:\Gamma_{E}\rightarrow G^{0}(R^{\mathrm{ps}}_{\overline{\Psi}})$ such that $\Psi^{u}=\Theta_{\psi^{u}}$. In particular,
(63) $\psi^{u}(\gamma)=1,\quad\forall\gamma\in[\Gamma_{E},\Gamma_{E}].$
Since $\Theta^{u}|_{\Gamma_{E}}$ is a deformation of $\overline{\Psi}$ to $R^{\mathrm{ps}}_{\overline{\Theta}}$ there is a homomorphism $R^{\mathrm{ps}}_{\overline{\Psi}}\rightarrow R^{\mathrm{ps}}_{\overline{\Theta}}$ such that
(64) $\Theta^{u}|_{\Gamma_{E}}=\Psi^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Psi}}}R^{\mathrm{ps}}_{\overline{\Theta}}.$
If $\rho\in X^{\mathrm{gen}}_{G,\overline{\Theta}}(A)$ then $\Theta_{\rho}=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}A$, and it follows from (64) that $\rho|_{\Gamma_{E}}=\psi^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Psi}}}A$. The assertion follows from (63). ∎
Theorem 9.3.
There is a finite extension $L^{\prime}$ of $L$ with the ring of integers $\mathcal{O}^{\prime}$, such that the following hold:
  1. (1)
    $G^{0}_{\mathcal{O}^{\prime}}$ is split and $(G/G^{0})_{\mathcal{O}^{\prime}}$ is a constant group scheme;
  2. (2)
    $\operatorname{Rep}^{\Gamma_{F}}_{G,\pi}(\mathcal{O}^{\prime})$ is non-empty.
Moreover, if $(1)$ and $(2)$ hold then there are isomorphisms of $\mathcal{O}^{\prime}$-algebras:
  • (3)
    $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}\cong\mathcal{O}^{\prime}\llbracket(\Gamma_{E}^{\mathrm{ab},p}\otimes M)^{\Delta}\rrbracket\cong\mathcal{O}^{\prime}[(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}]\llbracket x_{1},\ldots,x_{r}\rrbracket$;
  • (4)
    $A^{\mathrm{gen}}_{G,\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}\cong R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$;
  • (5)
    $R^{\square}_{\overline{\rho}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}\cong(R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime})\llbracket t_{1},\ldots,t_{s}\rrbracket\cong\mathcal{O}^{\prime}[(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}]\llbracket z_{1},\ldots,z_{r+s}\rrbracket$;
where $M$ is the character lattice of $G^{0}_{\mathcal{O}^{\prime}}$, $r=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot[F:\mathbb{Q}_{p}]+\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$, $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
Proof.
If $L^{\prime}$ is a finite extension of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and residue field $k^{\prime}$ then the functor $D^{\mathrm{ps}}_{\overline{\Theta}_{k^{\prime}}}:\mathfrak{A}_{\mathcal{O}^{\prime}}\rightarrow\operatorname{Set}$ is pro-representable by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$ and analogous statements hold for $X^{\mathrm{gen}}_{G,\overline{\Theta}_{k^{\prime}}}$ and $D^{\square}_{\overline{\rho}_{k^{\prime}}}$. Thus it is enough to prove the statement for $\mathcal{O}^{\prime}=\mathcal{O}$, while allowing to replace $L$ by a finite extension.
Since $G$ is a generalised torus, after replacing $L$ by a finite unramified extension we may assume that $G^{0}$ is split and $G/G^{0}$ is a constant group scheme. We may assume that $G/G^{0}=\underline{\Delta}$ as replacing $G$ by the preimage of $\underline{\Delta}$ does not change the functors under consideration. The character lattice $M$ of $G^{0}$ does not change if we further replace $L$ by a finite extension. As explained at the beginning of Section 5 we have an action of $\Delta$ on $M$.
We will apply the results of previous sections with $\Gamma_{1}=W_{E/F}$ and $\Gamma_{2}=E^{\times}$. It is a fundamental result of Langlands proved in [10] (see also a nice exposition by Birkbeck [2, Proposition 2.0.3]) that there are natural isomorphisms:
(65) $H_{1}(W_{E/F},M)\cong H_{1}(E^{\times},M)^{\Delta}\cong(E^{\times}\otimes M)^{\Delta}.$
[…]
$0\rightarrow E^{\times}\rightarrow\mathcal{E}\rightarrow I_{\Delta}\rightarrow 0$
and the image of the extension class in $H^{2}(\Delta,E^{\times})$ under the isomorphisms
$\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},E^{\times})\cong\operatorname{Ext}^{2}_{\mathbb{Z}[\Delta]}(\mathbb{Z},E^{\times})\cong H^{2}(\Delta,E^{\times})$
is equal to $[u_{E/F}]$.
Let $N$ be the pro-$p$ completion of $(\mathcal{E}\otimes M)_{\Delta}$. It follows from Proposition 8.9 and Euler–Poincaré characteristic formula that
(66) $h^{1}_{\mathrm{cont}}(\Gamma_{F},\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma_{F},\operatorname{ad}\overline{\rho})-h^{2}_{\mathrm{cont}}(\Gamma_{F},\operatorname{ad}\overline{\rho})=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N-\dim G_{k}.$
Lemma 9.1 and (66) ensure that the assumptions of Proposition 7.6 are satisfied, and hence after replacing $L$ by a finite extension we may ensure that $\operatorname{Rep}^{\widehat{W}_{E/F}}_{G,\pi}(\mathcal{O})$ is non-empty by Corollary 7.7. In particular, $\operatorname{Rep}^{\Gamma_{F}}_{G,\pi}(\mathcal{O})$ and $\operatorname{Rep}^{W_{E/F}}_{G,\pi}(\mathcal{O})$ are also non-empty.
Lemma 6.6 implies that there is $\rho\in X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}(R^{\mathrm{ps}}_{\overline{\Theta}})$ such that $\Theta^{u}=\Theta_{\rho}$. Since $\rho$ factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$ by Lemma 9.2 we conclude that $\Theta^{u}$ is obtained from a $G$-pseudocharacter of $\widehat{W}_{E/F}$ via inflation to $\Gamma_{F}$. This together with Lemmas 9.1 and 9.2 imply that in the definitions of the functors $D^{\mathrm{ps}}_{\overline{\Theta}}$, $D^{\square}_{\overline{\rho}}$ and $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ we can replace $\Gamma_{F}$ with $\widehat{W}_{E/F}$ without changing the functors themselves.
Lemma 7.10 and (65) imply that $R^{\mathrm{ps}}_{\overline{\Theta}}$ is isomorphic to $\mathcal{O}\llbracket N\rrbracket$, where $N$ is the pro-$p$ completion of $(E^{\times}\otimes M)^{\Delta}$. It follows from Lemma 8.7 that $N\cong(\Gamma^{\mathrm{ab}}_{E}\otimes M)^{\Delta}$, which is a finitely generated $\mathbb{Z}_{p}$-module of rank $r$ and torsion subgroup isomorphic to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$ by 8.8. This yields the isomorphisms in part (3). Part (4) follows from Proposition 7.13. Part (5) follows from part (4) and Lemma 7.4 as $E^{\times}/(E^{\times})^{p}$ is finite and the map $\rho\mapsto\Theta_{\rho}$ induces a map of local $\mathcal{O}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$. ∎
We will now deduce some corollaries, which hold without extending the scalars.
Corollary 9.4.
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ is formally smooth. In particular, it is flat and induces a bijection between the sets of irreducible components.
Proof.
Since the map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\mathrm{ps}}_{\overline{\Theta}}$ is faithfully flat, the assertion follows from [14, Tag 06CM] and part (5) of Theorem 9.3. ∎
Corollary 9.5.
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is smooth and induces a bijection between the sets of irreducible components. Moreover, $A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is flat over $\mathcal{O}$ of relative dimension $\dim G_{k}\cdot([F:\mathbb{Q}_{p}]+1)$.
Proof.
The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is formally smooth by the same argument as in Corollary 9.4 using part (4) of Theorem 9.3. Since it is of finite type it is smooth by [14, Tag 00TN]. The map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}:=\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is flat and of finite presentation and hence open by [14, Tag 00I1]. By [14, Tag 004Z] it is enough to show the fibres of $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ are irreducible.
To ease the notation we let $R=R^{\mathrm{ps}}_{\overline{\Theta}}$, $A=A^{\mathrm{gen}}_{G,\overline{\Theta}}$, $R^{\prime}=R\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$ and $A^{\prime}=A\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$. Let $x:\operatorname{Spec}\kappa\rightarrow\operatorname{Spec}R$ be a geometric point and let $x^{\prime}:\operatorname{Spec}\kappa\rightarrow\operatorname{Spec}R^{\prime}$ be a point above $x$. The map
$A\otimes_{R,x}\kappa\rightarrow A^{\prime}\otimes_{R^{\prime},x^{\prime}}\kappa$
is an isomorphism. Part (2) of Theorem 9.3 implies that
$A^{\prime}\otimes_{R^{\prime},x^{\prime}}\kappa\cong\kappa[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$
Thus the fibres of $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ are irreducible and the result follows.
The last assertion follows from the fact that $\mathcal{O}^{\prime}$ is finite and free over $\mathcal{O}$, parts (3) and (4) of Theorem 9.3 and $\dim G_{k}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M$. ∎
Lemma 9.6.
Let $A\in R^{\mathrm{ps}}_{\overline{\Theta}}\text{-}\mathrm{alg}$, let $\rho\in X^{\mathrm{gen}}_{G,\overline{\Theta}}(A)$ and let $\tau:G\hookrightarrow\mathbb{A}^{n}$ be a closed immersion of $\mathcal{O}$-schemes. Then $\tau(\rho(\Gamma_{F}))$ is contained in a finitely generated $R^{\mathrm{ps}}_{\overline{\Theta}}$-submodule of $A^{n}=\mathbb{A}^{n}(A)$.
Proof.
It is enough to verify the assertion after extending the scalars to $\mathcal{O}^{\prime}$ given by Theorem 9.3. This follows from Lemma 9.2 and Lemma 9.6. ∎
Remark 9.7.
Lemma 9.6 implies that the scheme $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ coincides with the scheme $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ defined in [11] for a generalised reductive group $G$, see [11, Proposition 7.3].
Corollary 9.8.
Let $p^{m}$ be the order of $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$. Assume that $\mathcal{O}$ contains all the $p^{m}$-th roots of unity and (1) and (2) in Theorem 9.3 hold with $\mathcal{O}^{\prime}=\mathcal{O}$. Then there is a canonical action of the character group $\mathrm{X}((\mu_{p^{\infty}}(E)\otimes M)^{\Delta})$ on the set of irreducible components of $\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$, $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ and $X^{\mathrm{gen}}_{G,\overline{\Theta}}$, respectively. Moreover, the following hold:
  1. (1)
    this action is faithful and transitive;
  2. (2)
    irreducible components of $\operatorname{Spec}R^{\square}_{\overline{\rho}}$ and $\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ are formally smooth over $\mathcal{O}$;
  3. (3)
    irreducible components of $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ are of the form
    $\operatorname{Spec}\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
    where $r$ and $s$ are as in Theorem 9.3.
Proof.
The assertion follows from 7.20. ∎
Corollary 9.9.
Let $\varphi:G\to H$ be a surjection of generalised tori over $\mathcal{O}$ and let $\overline{\rho}:\Gamma_{F}\to G(k)$ be a continuous representation. Then the map $(R^{\square}_{\varphi\circ\overline{\rho}}/\varpi)^{\mathrm{red}}\rightarrow(R^{\square}_{\overline{\rho}}/\varpi)^{\mathrm{red}}$ is flat, where $\mathrm{red}$ indicates reduced rings, and the fibre at the closed point has dimension $([F:\mathbb{Q}_{p}]+1)(\dim G_{k}-\dim H_{k})$.
Proof.
If $\mathcal{P}$ is a finitely generated $\mathbb{Z}_{p}$-module and $R=k\llbracket\mathcal{P}\rrbracket$ is the completed group algebra of $\mathcal{P}$ then $R^{\mathrm{red}}\cong k\llbracket\mathcal{P}^{\mathrm{tf}}\rrbracket$, where $\mathcal{P}^{\mathrm{tf}}$ is the maximal torsion free quotient of $\mathcal{P}$. In particular, $R^{\mathrm{red}}\cong k\llbracket x_{1},\ldots,x_{d}\rrbracket$, where $d=\dim_{\mathbb{Q}_{p}}(\mathcal{P}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$.
Let $f:\mathcal{P}_{1}\rightarrow\mathcal{P}_{2}$ be a homomorphism of finitely generated $\mathbb{Z}_{p}$-modules and let $R_{i}=k\llbracket\mathcal{P}_{i}\rrbracket$. If $\ker f$ is torsion then $\mathcal{P}_{1}^{\mathrm{tf}}$ is a submodule of $\mathcal{P}_{2}^{\mathrm{tf}}$ and the fibre $k\otimes_{R_{1}^{\mathrm{red}}}R_{2}^{\mathrm{red}}$ is isomorphic to $k\llbracket\mathcal{P}^{\mathrm{tf}}_{2}/\mathcal{P}^{\mathrm{tf}}_{1}\rrbracket$. Since the dimension of the fibre is equal to $d_{2}-d_{1}$, where $d_{i}=\dim_{\mathbb{Q}_{p}}(\mathcal{P}_{i}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$, we deduce from [14, Tag 00R4] that $R_{1}^{\mathrm{red}}\rightarrow R_{2}^{\mathrm{red}}$ is flat.
Let $M$ be the character lattice of $G^{0}$ and let $N$ be the character lattice of $H^{0}$. Then $N\subseteq M$ and it follows from (61) and 8.1 that the kernel of the map $((\mathcal{E}\otimes N)^{\wedge,p})_{\Delta}\rightarrow((\mathcal{E}\otimes M)^{\wedge,p})_{\Delta}$ is torsion. Since $R_{\overline{\rho}}^{\square}/\varpi\cong k\llbracket((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}\rrbracket$ by Lemma 7.5 and we may swap pro-$p$ completion with $\Delta$-coinvariants, we obtain the assertion by letting $\mathcal{P}_{1}=((\mathcal{E}\otimes N)_{\Delta})^{\wedge,p}$ and $\mathcal{P}_{2}=((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$. The assertion about the dimension of the fibre follows from 8.9. ∎

### v2

@@
 Let $L$ be another finite extension of $\mathbb{Q}_{p}$ with ring of integers $\mathcal{O}$, uniformiser $\varpi$ and residue field $k=\mathcal{O}/\varpi$.
-Let $G$ be a smooth affine group scheme over $\mathcal{O}$, such that its neutral component $G^{0}$ is torus and the component group $G/G^{0}$ is a finite étale group scheme over $\mathcal{O}$.
+Let $G$ be a smooth affine group scheme over $\mathcal{O}$, such that its neutral component $G^{0}$ is a torus and the component group $G/G^{0}$ is a finite étale group scheme over $\mathcal{O}$.
 We call such group schemes generalised tori.
@@
 Theorem 1.1.
-There is a finite extension $L^{\prime}$ of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and a continuous representation $\rho:\Gamma_{F}\rightarrow G(\mathcal{O}^{\prime})$ lifting $\overline{\rho}$.
+There is a finite extension $L^{\prime}$ of $L$ with ring of integers $\mathcal{O}^{\prime}$ and a continuous representation $\rho:\Gamma_{F}\rightarrow G(\mathcal{O}^{\prime})$ lifting $\overline{\rho}$.
 Remark 1.2.
-If $G\cong G^{0}\rtimes(G/G^{0})$ then a lift in Theorem 1.1 can be constructed over $\mathcal{O}$ using the Teichmüller lift to define a section of $G^{0}(\mathcal{O})\twoheadrightarrow G^{0}(k)$, which then induces a section $\sigma:G(k)\rightarrow G(\mathcal{O})$.
+If $G\cong G^{0}\rtimes(G/G^{0})$ then a lift as in Theorem 1.1 can be constructed over $\mathcal{O}$ using the Teichmüller lift to define a section of $G^{0}(\mathcal{O})\twoheadrightarrow G^{0}(k)$, which then induces a section $\sigma:G(k)\rightarrow G(\mathcal{O})$.
 Then $\sigma\circ\overline{\rho}$ is the required lift.
@@
 Theorem 1.3.
-Assume that Theorem 1.1 holds with $\mathcal{O}=\mathcal{O}^{\prime}$ then
+Assume that Theorem 1.1 holds with $\mathcal{O}=\mathcal{O}^{\prime}$. Then
 $R^{\square}_{\overline{\rho}}\cong\mathcal{O}[(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}]\llbracket x_{1},\ldots,x_{m}\rrbracket,$
@@
 If we let $G^{\prime}$ be the derived subgroup scheme of $G^{0}$, then $G^{0}/G^{\prime}$ is a torus and $G/G^{\prime}$ is a generalised torus as considered in this paper.
-If $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ is a continuous representation then we show in [11, Proposition 4.20] that its deformation ring $R^{\square}_{\overline{\rho}}$ can be presented over the deformation ring of $\varphi\circ\overline{\rho}:\Gamma_{F}\rightarrow(G/G^{\prime})(k)$, where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map.
+If $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ is a continuous representation then we show in [12, Proposition 4.20] that its deformation ring $R^{\square}_{\overline{\rho}}$ can be presented over the deformation ring of $\varphi\circ\overline{\rho}:\Gamma_{F}\rightarrow(G/G^{\prime})(k)$, where $\varphi:G\rightarrow G/G^{\prime}$ is the quotient map.
 This allows us to split up the arguments into “torus part”, carried out in this paper, and “semisimple part” carried out in [12].
@@
 Further, one may interpret $\widehat{Z}^{1}$ as $D^{\square}_{\overline{\rho}}$ for the group $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ given by $\overline{\rho}(\gamma)=(1,\pi(\gamma))$, where $\pi:\Gamma_{F}\rightarrow\Delta$ is the quotient map.
-So that $\widehat{Z}^{1}$ is the analogue of $D^{\square}_{\mathbf{1}}$ and we have to find the ring pro-representing the functor $\widehat{Z}^{1}$.
+The functor $\widehat{Z}^{1}$ is the analogue of $D^{\square}_{\mathbf{1}}$ and we have to find the ring pro-representing the functor $\widehat{Z}^{1}$.
 Before we describe our solution, let us point out one dead end. If $\Phi\in\widehat{Z}^{1}(A)$ then its restriction to $\Gamma_{E}$ is just a continuous homomorphism
@@
 We conclude that the restriction to $\Gamma_{E}$ will lose information in general.
-Our solution is motivated by Langlands paper [10] on his correspondence for tori, where he shows that the universal coefficient theorem induces an isomorphism
+Our solution is motivated by Langlands’ paper [11] on his correspondence for tori, where he shows that the universal coefficient theorem induces an isomorphism
 (1) $H^{1}(W_{E/F},\operatorname{Hom}(M,\mathbb{C}^{\times}))\cong\operatorname{Hom}(H_{1}(W_{E/F},M),\mathbb{C}^{\times}))$
@@
 corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$.
-Since every continuous representation $\Gamma_{F}\rightarrow G(A)$ factors through the profinite completion of $W_{E/F}$, this motivates us to study the functor
+Since every continuous representation $\Gamma_{F}\rightarrow G(A)$ with image in $(G/G^{0})(A)$ equal to $\Delta$ factors through the profinite completion of $W_{E/F}$, this motivates us to study the functor
 $\operatorname{CRing}\rightarrow\operatorname{Ab},\quad A\mapsto Z^{1}(W_{E/F},G(A))$
@@
 $\operatorname{CRing}\rightarrow\operatorname{Ab},\quad A\mapsto H^{1}(\Gamma_{1},\operatorname{Hom}(M,A^{\times}))$
-is representable by the group algebra $\mathbb{Z}[H_{1}(\Gamma_{1},M)]$.
+is represented by the group algebra $\mathbb{Z}[H_{1}(\Gamma_{1},M)]$.
 We prove 1.7 by first replacing $M$ with $\mathbb{Z}[\Delta]\otimes M$. In this case
@@
 To prove 1.7 we verify that the action of $\Delta$ on $(\Gamma_{2}^{\mathrm{ab}}\oplus I_{\Delta})\otimes M$ induced by its action on the functor in (6) is isomorphic to $\mathcal{E}\otimes M$.
-This calculation is carried out in Sections 4.
+This calculation is carried out in Section 4.
 The last part of Theorem 1.7 follows by observing that (1) remains an isomorphism if $\mathbb{C}^{\times}$ is replaced by any divisible group and for any $A$ there is a faithfully flat map $A\rightarrow B$ such that $B^{\times}$ is divisible.
@@
 Mapping a representation to its $G$-pseudocharacter induces an isomorphism of schemes
-$\operatorname{Rep}^{\Gamma_{1}}_{G}\sslash G^{0}\overset{\cong}{\longrightarrow}\mathrm{PC}^{\Gamma_{1}}_{G}$
+$\operatorname{Rep}^{\Gamma_{1}}_{G}\sslash G^{0}\overset{\cong}{\longrightarrow}\mathrm{PC}^{\Gamma_{1}}_{G}.$
 The proof of 1.8 follows the arguments of Emerson–Morel [8], where they prove an analogous result, when $G$ is a connected reductive group over a field of characteristic 00.
@@
 Let us assume that $\operatorname{Rep}^{\Gamma_{1}}_{G}(\mathcal{O})$ is non-empty, $G^{0}$ is split over $\mathcal{O}$ and $G/G^{0}$ is a constant group scheme.
-Let $\rho_{0}:\Gamma_{1}\rightarrow G(\mathcal{O})$ be a representation and let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ be the subfunctor of $\operatorname{Rep}^{\Gamma_{1}}_{G}$ consisting of representations $\rho$ such that $\Pi\circ\rho=\Pi\circ\rho_{0}$, where $\Pi:G\rightarrow G/G^{0}$.
+Let $\rho_{0}:\Gamma_{1}\rightarrow G(\mathcal{O})$ be a representation and let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ be the subfunctor of $\operatorname{Rep}^{\Gamma_{1}}_{G}$ consisting of representations $\rho$ such that $\Pi\circ\rho=\Pi\circ\rho_{0}$, where $\Pi:G\rightarrow G/G^{0}$ is the quotient map.
 One may similarly define $\mathrm{PC}^{\Gamma_{1}}_{G,\pi}$.
@@
 1.3. Producing a lift
-Let us go back to the problem of exhibiting a lift of $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ to the characteristic zero as claimed in Theorem 1.1.
+Let us go back to the problem of exhibiting a lift of $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ to characteristic zero as claimed in 1.1.
 Using $\overline{\rho}$ instead of $\rho_{0}$, we may identify the restriction of $D^{\square}_{\overline{\rho}}$ to $\mathfrak{A}_{k}$ with the restriction of $\widehat{Z}^{1}$ to $\mathfrak{A}_{k}$.
@@
 In Section 8 we compute
-$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}(\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot([F:\mathbb{Q}_{p}]+1),$
+$\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M\cdot([F:\mathbb{Q}_{p}]+1),$
 which is then also equal to $\dim R^{\square}_{\overline{\rho}}/\varpi$. On the other hand, using Mazur’s obstruction theory and the Euler–Poincaré characteristic formula we show that
@@
 If $\mathcal{A}$ is an abelian group then we will denote its pro-$p$ completion by $\mathcal{A}^{\wedge,p}$.
-We denote by $\operatorname{Set}$, $\operatorname{Ab}$, $\operatorname{CRing}$ the categories of sets, abelian groups and commutative rings, respectively.
-If $X$ is a scheme the we denote the ring of its global sections by $\mathcal{O}(X)$. If $\Gamma$ is a group then we denote by $\mathcal{O}[\Gamma]$ its group algebra over a ring $\mathcal{O}$.
+If $X$ is a scheme then we denote the ring of its global sections by $\mathcal{O}(X)$. If $\Gamma$ is a group then we denote by $\mathcal{O}[\Gamma]$ its group algebra over a ring $\mathcal{O}$.
 1.6. Acknowledgements
@@
 The authors would like to thank Toby Gee and James Newton for their comments on a draft version of this paper.
+The authors thank the anonymous referee for their helpful comments that improved the quality of the manuscript.
 Parts of the paper were written during the research stay of VP at the Hausdorff Research Institute for Mathematics in Bonn for the Trimester Program The Arithmetic of the Langlands Program.
@@
 Remark 2.2.
-It follows from [5, Proposition 3.1.3], that for a generalised torus $G$ the quotient $G/G^{0}$ is étale.
+It follows from [6, Proposition 3.1.3], that for a generalised torus $G$ the quotient $G/G^{0}$ is étale.
 Sean Cotner has shown in [7] that if $S$ is locally noetherian then for a smooth affine $S$-group scheme $G$ the identity component $G^{0}$ has reductive geometric fibers and $G/G^{0}$ is finite étale if and only if $G$ is geometrically reductive in the sense of [1, Definition 9.1.1].
@@
 ${(M\otimes_{A}N)^{G^{0}}=M^{G^{0}}\otimes_{A}N}.$
-If $M$ is an abelian group then we denote by $\mathfrak{D}(M):=\operatorname{Spec}(\mathcal{O}[M])$ the diagonalizable group scheme, where the Hopf algebra structure on $\mathcal{O}[M]$ is given by addition in $M$ and inversion.
+If $M$ is an abelian group then we denote by $\mathfrak{D}(M):=\operatorname{Spec}(\mathcal{O}[M])$ the diagonalisable group scheme, where the Hopf algebra structure on $\mathcal{O}[M]$ is given by addition and inversion in $M$.
 For any $\mathcal{O}$-algebra $A$, we have
@@
 Proof.
-[9, Theorem 3.7]. ∎
+[10, Theorem 3.7]. ∎
 We will construct an explicit inverse of the map above following [10].
@@
 Lemma 3.2.
-Let $\varphi:\Gamma_{2}\rightarrow V$ be a $1$-cocycle, and let $f_{\varphi}:\Gamma_{2}\rightarrow\mathcal{F}_{V}$ be a $1$-cocycle corresponding to $\varphi$, via the canonical isomorphism $\mathcal{F}_{V}\cong V$ and let ${\Phi_{\varphi}:\Gamma_{1}\rightarrow\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}}$ be a function given by
+Let $\varphi:\Gamma_{2}\rightarrow V$ be a $1$-cocycle, and let $f_{\varphi}:\Gamma_{2}\rightarrow\mathcal{F}_{V}$ be the $1$-cocycle corresponding to $\varphi$, via the canonical isomorphism $\mathcal{F}_{V}\cong V$ and let ${\Phi_{\varphi}:\Gamma_{1}\rightarrow\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{V}}$ be the function given by
 $\Phi_{\varphi}(g)=\sum_{c}\overline{c}^{-1}f_{\varphi}(\overline{c}g\overline{cg}^{-1}).$
@@
 However once these representatives have been fixed it is immediate from the formulas that the map is functorial in $V$.
-So that if $\alpha:V\rightarrow W$ is a $\Gamma_{2}$-equivariant homomorphism of abelian groups equipped with $\Gamma_{2}$-action and $\psi\in Z^{1}(\Gamma_{2},W)$ is the image of $\varphi$ under the map $Z^{1}(\Gamma_{2},V)\rightarrow Z^{1}(\Gamma_{2},W)$ induced by $\alpha$ then $\Phi_{\varphi}$ maps to $\Phi_{\psi}\in Z^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{W})$.
+If $\alpha:V\rightarrow W$ is a $\Gamma_{2}$-equivariant homomorphism of abelian groups equipped with $\Gamma_{2}$-action and $\psi\in Z^{1}(\Gamma_{2},W)$ is the image of $\varphi$ under the map $Z^{1}(\Gamma_{2},V)\rightarrow Z^{1}(\Gamma_{2},W)$ induced by $\alpha$ then $\Phi_{\varphi}$ maps to $\Phi_{\psi}\in Z^{1}(\Gamma_{1},\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}{W})$.
 4. The space of $1$-cocycles
@@
 Let $V$ be an abelian group equipped with a trivial $\Gamma_{2}$-action.
-We may identify $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ with the space of functions $f:\Delta\rightarrow V$ and $[g\cdot f](c)=f(cg)$ for all $g\in\Gamma_{1}$ and $c\in\Delta$.
+We may identify $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ with the space of functions $f:\Delta\rightarrow V$ where the action is given by $[g\cdot f](c)=f(cg)$ for all $g\in\Gamma_{1}$ and $c\in\Delta$.
 We also have an action of $\Delta$ on $\operatorname{Ind}_{\Gamma_{2}}^{\Gamma_{1}}V$, which commutes with the action of $\Gamma_{1}$ and is given by $[d\cdot f](c)=f(d^{-1}c)$ for all $d\in\Delta$.
@@
 We want to understand the action of $\Delta$ on $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V)\oplus\operatorname{Hom}(\Gamma_{2},V)$ via this isomorphism.
-It follows from (8) that $[[d\ast\Phi](g)](cd)=[\Phi(g)](c)$.
+It follows from (8) that $[[d\ast\Phi](g)](dc)=[\Phi(g)](c)$.
 Hence
-$[d\ast\Phi_{\varphi}](g)=\sum_{c\in\Delta}\varphi(\overline{c}g\overline{cg}^{-1})\mathbf{1}_{cd}.$
+$[d\ast\Phi_{\varphi}](g)=\sum_{c\in\Delta}\varphi(\overline{c}g\overline{cg}^{-1})\mathbf{1}_{dc}.$
 for $\varphi\in\operatorname{Hom}(\Gamma_{2},V)$.
@@
 Hence, $\overline{ck}=\overline{c}$ and we conclude that
-$[[d\ast\Phi_{\varphi}](k)](1)=\sum_{c}\varphi(\overline{c}k\overline{c}^{-1})\mathbf{1}_{cd}(1)=\varphi(\overline{d^{-1}}k\overline{d^{-1}}^{-1})=\varphi(\overline{d}^{-1}k\overline{d})=[d\cdot\varphi](k).$
-We note that $\varphi((\overline{d}h)^{-1}k(\overline{d}h))=\varphi(\overline{d}^{-1}k\overline{d})$ for all $h,k\in\Gamma_{2}$, which justifies the last two equalities.
-Thus the difference $d\ast\Phi_{\varphi}-\Phi_{d\cdot\varphi}$ is a $1$-coboundary and we want to find a function $f:\Delta\rightarrow V$ depending on $d$ such that for all $g\in\Gamma_{1}$ we have
+$[[d\ast\Phi_{\varphi}](k)](1)=\sum_{c}\varphi(\overline{c}k\overline{c}^{-1})\mathbf{1}_{dc}(1)=\varphi(\overline{d^{-1}}k\overline{d^{-1}}^{-1})=\varphi(\overline{d}^{-1}k\overline{d})=:[d\cdot\varphi](k).$
+We note that $\varphi((\overline{d}h)^{-1}k(\overline{d}h))=\varphi(\overline{d}^{-1}k\overline{d})$ for all $h,k\in\Gamma_{2}$, which justifies the third equality and we take the last equality as the definition $d\cdot\varphi$.
+Since $[[\Phi_{d\cdot\varphi}](k)](1)=[d\cdot\varphi](k)$ by 3.2, 3.1 implies that the cocycle $d\ast\Phi_{\varphi}-\Phi_{d\cdot\varphi}$ is a $1$-coboundary and thus there is a function $f:\Delta\rightarrow V$ depending on $d$ such that for all $g\in\Gamma_{1}$ we have
 (11) $[d\ast\Phi_{\varphi}](g)-\Phi_{d\cdot\varphi}(g)=(g-1)f.$
@@
 For $d,c\in\Delta$ we define $\kappa(d,c)$ to be the image of $\overline{d}\overline{c}\overline{dc}^{-1}$ in $\Gamma_{2}^{\mathrm{ab}}$.
-Then $\kappa:\Delta\times\Delta\rightarrow\Gamma_{2}^{\mathrm{ab}}$ is the $2$-cocycle corresponding to the extension
+The map $\kappa:\Delta\times\Delta\rightarrow\Gamma_{2}^{\mathrm{ab}}$ is the $2$-cocycle corresponding to the extension
 $0\rightarrow\Gamma_{2}^{\mathrm{ab}}\rightarrow\Gamma_{1}/[\Gamma_{2},\Gamma_{2}]\rightarrow\Delta\rightarrow 0$
@@
 $\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})\cong\operatorname{Ext}^{2}_{\mathbb{Z}[\Delta]}(\mathbb{Z},\Gamma_{2}^{\mathrm{ab}})\cong H^{2}(\Delta,\Gamma_{2}^{\mathrm{ab}}),$
-see [3, Chapter XIV, §4, Remark].
+see [4, Chapter XIV, §4, Remark].
 The extension class of (17) gives an element in $\operatorname{Ext}^{1}_{\mathbb{Z}[\Delta]}(I_{\Delta},\Gamma_{2}^{\mathrm{ab}})$.
@@
 (19) $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)\cong\operatorname{Hom}(\mathbb{Z}[\Delta],\operatorname{Hom}(M,V))\cong\operatorname{Hom}(\mathbb{Z}[\Delta]\otimes M,V),$
-where the action of $\Gamma_{1}$ on the last term is induced its action on $\mathbb{Z}[\Delta]\otimes M$, which is given by $g\cdot(c\otimes m)=cg^{-1}\otimes m$.
-The diagonal action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$ is given by $d\cdot(c\otimes m)=dc\otimes dm$, commutes with the action of $\Gamma_{1}$ and induces an action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)$, which commutes with the action of $\Gamma_{1}$.
+where the action of $\Gamma_{1}$ on the last term is induced by its action on $\mathbb{Z}[\Delta]\otimes M$, which is given by $g\cdot(c\otimes m)=cg^{-1}\otimes m$, so that $[g\cdot f](x)=f(g^{-1}x)$ for all $x\in\mathbb{Z}[\Delta]\otimes M$.
+The diagonal action of $\Delta$ on $\mathbb{Z}[\Delta]\otimes M$ is given by $d\cdot(c\otimes m)=dc\otimes dm$, commutes with the action of $\Gamma_{1}$ and induces a (left) action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V)$, which commutes with the action of $\Gamma_{1}$.
 If $M=\mathbb{Z}$ with the trivial $\Delta$-action then the construction recovers the action of $\Delta$ on $\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}V$ considered earlier.
 Lemma 4.4.
-$(\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))^{\Delta}$ as $\Gamma_{1}$-representation is naturally isomorphic to $\operatorname{Hom}(M,V)$ with $\Gamma_{1}$ acting on it via $\Delta$.
+As $\Gamma_{1}$-representations $(\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))^{\Delta}$ is naturally isomorphic to $\operatorname{Hom}(M,V)$ with $\Gamma_{1}$ acting on it via $\Delta$.
 Proof.
@@
 $[\vartheta(\alpha)(c)](m)=\alpha(c^{-1}m),\quad\forall c\in\Delta,\quad\forall m\in M.$
-Let $b\in B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ be a boundary $b(g)=(g-1)\alpha$.
+Let $b\in B^{1}(\Gamma_{1},\operatorname{Hom}(M,V))$ be a boundary with $b(g)=(g-1)\alpha$ for all $g\in\Gamma_{1}$.
 Its image in $B^{1}(\Gamma_{1},\operatorname{Ind}^{\Gamma_{1}}_{\Gamma_{2}}\operatorname{Hom}(M,V))$ is the boundary $b^{\prime}$ given by $b^{\prime}(g)=(g-1)\vartheta(\alpha)$.
@@
 Lemma 4.7.
-If $V$ is any abelian group then for all $\Gamma_{1}$-modules $N$ we have a canonical map
-(21) $H^{1}(\Gamma_{1},\operatorname{Hom}(N,V))\rightarrow\operatorname{Hom}(H_{1}(\Gamma_{1},N),V),$
+If $V$ is any abelian group then for all $\Gamma_{1}$-modules $N$ and all $i\geq 0$ we have a canonical map
+(21) $H^{i}(\Gamma_{1},\operatorname{Hom}(N,V))\rightarrow\operatorname{Hom}(H_{i}(\Gamma_{1},N),V),$
 which is an isomorphism if $V$ is divisible.
@@
 Lemmas 4.6 and 4.8 give us the commutative diagram above.
-Moreover, if $V$ is divisible then the top row is exact, the last arrow in the top row is surjective and the first vertical arrow is an isomorphism.
+Moreover, if $V$ is divisible then by 4.7 the top row is exact, the last arrow in the top row is surjective and the first vertical arrow is an isomorphism.
 We deduce that for all divisible $V$ the maps $q:(\mathcal{E}\otimes M)_{\Delta}\twoheadrightarrow I_{\Delta}M$ and $\varphi_{\mathrm{nat}}:H_{1}(\Gamma,M)\rightarrow(\mathcal{E}\otimes M)_{\Delta}$ induce an exact sequence
@@
 If $M$ is a free $\mathbb{Z}$-module of finite rank then
-(25) $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]\cong\mathbb{Z}[H_{1}(\Gamma,M)][t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
+(25) $\mathbb{Z}[(\mathcal{E}\otimes M)_{\Delta}]\cong\mathbb{Z}[H_{1}(\Gamma_{1},M)][t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}],$
 where $s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M-\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M_{\Delta}$.
@@
 Proof.
-The right hand side is an fpqc sheaf, since it is representable [14, Tag 023Q].
+The right hand side is an fpqc sheaf, since it is representable [15, Tag 023Q].
 So it suffices to find for every ring $A$ and every homomorphism ${f:\mathbb{Z}[H_{1}(\Gamma_{1},M)]\to A}$ an fpqc cover of $A$ and a descent datum in the left hand side, which maps to a descent datum in the right hand side, which descends to $f$.
-By [16, Lemma 4.1.1] for every ring $A$ there is a faithfully flat map $A\to B$, such that $B^{\times}$ is divisible.
+By [17, Lemma 4.1.1] for every ring $A$ there is a faithfully flat map $A\to B$, such that $B^{\times}$ is divisible.
 By 4.7 the map
@@
 We say that a representation $\rho:\Gamma_{1}\rightarrow G(A)$ is admissible if $\Pi\circ\rho:\Gamma_{1}\to\underline{\Delta}(A)$, sends $\gamma\in\Gamma_{1}$ to the image of $\pi(\gamma)\in\Delta$ in $\underline{\Delta}(A)$.
-This terminology is motivated by admissible Galois representations into $L$-groups appearing in the Langlands correspondence.
+This terminology is motivated by admissible Galois representations into $L$-groups appearing in the Langlands correspondence, see [3, §9].
 Let $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ be the set of admissible representations $\rho:\Gamma_{1}\rightarrow G(A)$.
@@
 In particular, $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ is representable by an $\mathcal{O}$-algebra and (29) induces a natural isomorphism
-(30) $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\xrightarrow{\cong}\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$
+(30) $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\xrightarrow{\cong}\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}].$
 Moreover, (29) is $\mathfrak{D}(M)(A)$-equivariant and induces a natural bijection between the set of $\mathfrak{D}(M)(A)$-orbits in $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ and $H^{1}(\Gamma_{1},\mathfrak{D}(M)(A))$.
@@
 Corollary 5.5.
+Assume that there is a representation $\rho_{0}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$.
 The action of $\mathfrak{D}(M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ factors through the action of $\mathfrak{D}(I_{\Delta}M)$, which acts freely.
@@
 […]
+Let $A$ be any $\mathcal{O}$-algebra and let $H(A)$ be the group $\operatorname{Hom}((\mathcal{E}\otimes M)_{\Delta},A^{\times})$.
+Since $q:(\mathcal{E}\otimes M)_{\Delta}\rightarrow I_{\Delta}M$ is surjective, we may identify $\mathfrak{D}(I_{\Delta}M)(A)$ with a subgroup of $H(A)$.
+The isomorphism (30) identifies $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ with $H(A)$ and the action of $\mathfrak{D}(I_{\Delta}M)(A)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(A)$ is identified with the action of $\mathfrak{D}(I_{\Delta}M)(A)$ on $H(A)$ by multiplication, which is a free action.
+Hence, the action of $\mathfrak{D}(I_{\Delta}M)$ on $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}$ is free.
+∎
 Corollary 5.6.
@@
 Proof.
-Proposition 5.2 implies that it is enough to prove the statement, when $A=\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ and $\rho=\Phi_{\mathrm{nat}}\rho_{0}$.
+Proposition 5.2 implies that it is enough to prove the statement, when $A=\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}]$ and $\rho=\Phi_{\mathrm{nat}}\rho_{0}$, where $\rho_{0}$ is any representation in $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ and $\Phi_{\mathrm{nat}}$ is the cocycle defined in Section 4.
 Since $\Gamma_{2}$ is of finite index in $\Gamma_{1}$ it is enough to show that $\tau(\rho(\Gamma_{2}))$ is contained in a finitely generated $\mathcal{O}[H_{1}(\Gamma_{1},M)]$-submodule of $(\mathcal{O}[(\mathcal{E}\otimes M)_{\Delta}])^{n}$.
@@
 We keep the notation of Section 5.
-We now recall Lafforgue’s notion of $G$-pseudocharacter in the form of [12, Definition 3.1].
-The difference to Lafforgue’s original definition [8, Section 11] is that we work over the base ring $\mathcal{O}$ and allow $G$ to be disconnected.
+We now recall Lafforgue’s notion of $G$-pseudocharacter in the form of [13, Definition 3.1].
+The difference to Lafforgue’s original definition [9, Section 11] is that we work over the base ring $\mathcal{O}$ and allow $G$ to be disconnected.
 The definition works for generalised reductive $\mathcal{O}$-group schemes as defined in [12], but this generality is not needed here.
@@
 We denote the set of $G$-pseudocharacters of $\Gamma_{1}$ over $A$ by $\mathrm{PC}_{G}^{\Gamma_{1}}(A)$.
-The functor $A\mapsto\mathrm{PC}_{G}^{\Gamma_{1}}(A)$ is representable by an affine $\mathcal{O}$-scheme $\mathrm{PC}_{G}^{\Gamma_{1}}$ [12, Theorem 3.20].
+The functor $A\mapsto\mathrm{PC}_{G}^{\Gamma_{1}}(A)$ is representable by an affine $\mathcal{O}$-scheme $\mathrm{PC}_{G}^{\Gamma_{1}}$ [13, Theorem 3.20].
 When $\varphi:G\to H$ is a homomorphism of generalised tori over $\mathcal{O}$, the induced maps $\varphi^{*}_{n}:\mathcal{O}[H^{n}]^{H^{0}}\to\mathcal{O}[G^{n}]^{G^{0}}$ give rise to an $H$-pseudocharacter $(\Theta_{n}\circ\varphi^{*}_{n})_{n\geq 1}$.
 By analogy with the notation for representations we denote this $H$-pseudocharacter by $\varphi\circ\Theta$.
-So we also have an induced map $\mathrm{PC}_{G}^{\Gamma}(A)\to\mathrm{PC}_{H}^{\Gamma}(A)$.
+Thus we also have an induced map $\mathrm{PC}_{G}^{\Gamma}(A)\to\mathrm{PC}_{H}^{\Gamma}(A)$.
 It is easy to verify that specialisation along $f:A\to B$ commutes with composition with $\varphi$, i.e.
@@
 We will show, that (37) and (39) are isomorphisms.
-The result is specific to generalised tori [^2] We don’t expect (37) and (39) to be isomorphisms when $G$ is an arbitrary generalised reductive group in the sense of [12, Definition 2.3].
+The result is specific to generalised tori [^2] We don’t expect (37) and (39) to be isomorphisms when $G$ is an arbitrary generalised reductive group in the sense of [13, Definition 2.3].
 However, we don’t know of an example where (37) is not an isomorphism.
@@
 Proof.
-As $G^{0}$ is linearly reductive the proof of [7, Proposition 2.11 (i)] applies, we recall the argument here.
+As $G^{0}$ is linearly reductive the proof of [8, Proposition 2.11 (i)] applies, we recall the argument here.
 Let $\mathcal{O}(G^{\Gamma_{1}})$ be the $\mathcal{O}$-algebra which represents the functor
@@
 We also have a natural surjection $\mathcal{O}(G^{\Gamma_{1}})^{G^{0}}\twoheadrightarrow\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})$ with kernel $I$.
-The structure of $J$ and $I$ is described explicitly in [7, Proposition 2.5]; note that there $G$ is assumed to be connected, but the description generalises easily to our situation.
+The structure of $J$ and $I$ is described explicitly in [8, Proposition 2.5]; note that there $G$ is assumed to be connected, but the description generalises easily to our situation.
 Namely $J$ is generated by the image of the $G^{0}$-equivariant maps
@@
 $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})\otimes_{\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}\cong\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G})\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$
-by 6.3.
-We conclude, that (41) is obtained from (40) by tensoring with $-\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$.
-So by 6.2 (41) is an isomorphism as well.
-∎
+by 6.3. We conclude, that (41) is obtained from (40) by applying $-\otimes_{\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{\underline{\Delta}})}\mathcal{O}$. So by 6.2 (41) is an isomorphism as well. ∎
 Corollary 6.5.
@@
 Conversely, if $I$ is an ideal of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})_{\mathfrak{m}}$ such that the quotient $A$ is finite then $A\in\mathfrak{A}_{\mathcal{O}}$.
-By specialising the universal representation $\Gamma_{1}\rightarrow G(\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}))$ along $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow A$ we obtain a representation $\rho:\Gamma_{1}\rightarrow G(A)$.
+The map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow k$ corresponding to $\overline{\rho}$ factors through $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\rightarrow k$ as $\overline{\rho}|_{\Gamma_{1}}\in\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(k)$ by assumption.
+Since $G/G^{0}$ is constant and $A$ is a local ring we have $(G/G^{0})(A)=(G/G^{0})(k)$ and hence the map $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G})\rightarrow A$ factors through $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})\rightarrow A$.
+By specialising the universal representation $\Gamma_{1}\rightarrow G(\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}))$ along this map we obtain a representation $\rho:\Gamma_{1}\rightarrow G(A)$.
 Since the target is finite $\rho$ extends uniquely to a continuous representation $\tilde{\rho}:\widehat{\Gamma}_{1}\rightarrow G(A)$, which defines a point in $D^{\square}_{\overline{\rho}}(A)$.
@@
 The equivalence of (1) and (2) follows from Nakayama’s lemma and the fact that $N/pN$ is the pro-$p$ completion of $\mathcal{A}/p\mathcal{A}\cong\oplus_{I}\mathbb{F}_{p}$ for some set $I$.
-If $N$ is finitely generated as $\mathbb{Z}_{p}$-module, then $N\cong\mu\oplus\mathbb{Z}_{p}^{r}$, where $\mu$ is a finite $p$-group and $r=\dim_{\mathbb{Q}_{p}}N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$.
+If $N$ is finitely generated as a $\mathbb{Z}_{p}$-module, then $N\cong\mu\oplus\mathbb{Z}_{p}^{r}$, where $\mu$ is a finite $p$-group and $r=\dim_{\mathbb{Q}_{p}}N\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p}$.
 Hence, $\mathcal{O}\llbracket N\rrbracket\cong\mathcal{O}[\mu]\llbracket x_{1},\ldots,x_{r}\rrbracket$ and hence (1) implies (3).
@@
 It follows from (17) that if $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is finite then $\mathcal{E}/p\mathcal{E}$ is also finite.
-Since $M$ is a finite free $\mathbb{Z}$-module, the assertion follows from Lemma 7.3 applied to $\mathcal{A}=(\mathcal{E}\otimes M)_{\Delta}$.
+Since $M$ is a finite free $\mathbb{Z}$-module, the assertion follows from 7.1 and Lemma 7.3 applied to $\mathcal{A}=(\mathcal{E}\otimes M)_{\Delta}$.
 ∎
@@
 Moreover, if $h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})$ is also finite and
-(46) $h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\leq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}-\dim G_{k}$
+(46) $h^{1}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{0}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})-h^{2}_{\mathrm{cont}}(\Gamma,\operatorname{ad}\overline{\rho})\geq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}-\dim G_{k}$
 then $R^{\square}_{\overline{\rho}}$ is complete intersection, $\mathcal{O}$-flat of relative dimension $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$.
@@
 Let $N:=((\mathcal{E}\otimes M)_{\Delta})^{\wedge,p}$.
-By considering (47) modulo $\varpi$ and using Lemma 7.5 we deduce that $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N\leq r$.
-Moreover, the assumption $r-s\leq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ implies that $r-s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ and $\varpi,f_{1},\ldots,f_{s}$ can be extended to a system of parameters of a regular ring $\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket$.
+By considering (47) modulo $\varpi$ and using Lemma 7.5 we deduce that $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N\geq r-s$.
+Moreover, the assumption (46) implies that $r-s\geq\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$.
+Hence, $r-s=\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$ and $\varpi,f_{1},\ldots,f_{s}$ can be extended to a system of parameters of a regular ring $\mathcal{O}\llbracket x_{1},\ldots,x_{r}\rrbracket$.
 Thus $\varpi,f_{1},\ldots,f_{s}$ are a part of a regular sequence, and hence $R^{\square}_{\overline{\rho}}$ is complete intersection, flat over $\mathcal{O}$ of relative dimension of $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}N$.
@@
 Proof.
-As the image of $\Gamma_{1}$ in $\widehat{\Gamma}_{1}$ is dense, injectivity follows from [12, Lemma 3.2].
+As the image of $\Gamma_{1}$ in $\widehat{\Gamma}_{1}$ is dense, injectivity follows from [13, Lemma 3.2].
 Let $\Theta\in\mathrm{PC}^{\Gamma_{1}}_{G,\pi}(A)$.
@@
 Let $\overline{\Theta}$ be the $G$-pseudocharacter of $\widehat{\Gamma}_{1}$ associated to $\overline{\rho}$.
-Let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the deformation functor which sends $A$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\widehat{\Gamma}_{1}$ valued in $A$ such that $\Theta\otimes_{A}k=\overline{\Theta}.$ The functor $D^{\mathrm{ps}}_{\overline{\Theta}}$ is pro-represented by $R^{\mathrm{ps}}_{\overline{\Theta}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ by [12, Theorem 5.4].
+Let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the deformation functor which sends $A$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\widehat{\Gamma}_{1}$ valued in $A$ such that $\Theta\otimes_{A}k=\overline{\Theta}.$ The functor $D^{\mathrm{ps}}_{\overline{\Theta}}$ is pro-represented by $R^{\mathrm{ps}}_{\overline{\Theta}}\in\widehat{\mathfrak{A}}_{\mathcal{O}}$ by [13, Theorem 5.4].
 Proposition 7.9.
@@
 Proof.
-Since $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty Corollary 6.5 allows us to identify $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[H_{1}(\Gamma,M)]$.
+Since $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty Corollary 6.5 allows us to identify $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ with $\mathcal{O}[H_{1}(\Gamma,M)]$.
 The assertion follows from Proposition 7.9 and Lemma 7.2.
@@
 Lemma 7.11.
-If $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is a finite $p$-group and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian and is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}(\operatorname{Rep}^{\Gamma_{1}}_{G,\pi})$, where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ corresponding to $\overline{\Theta}|_{\Gamma_{1}}$.
+If $\Gamma_{2}^{\mathrm{ab}}/p\Gamma_{2}^{\mathrm{ab}}$ is a finite $p$-group and $\operatorname{Rep}^{\Gamma_{1}}_{G,\pi}(\mathcal{O})$ is non-empty then $R^{\mathrm{ps}}_{\overline{\Theta}}$ is noetherian and is naturally isomorphic to the $\mathfrak{m}$-adic completion of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$, where $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}(\mathrm{PC}^{\Gamma_{1}}_{G,\pi})$ corresponding to $\overline{\Theta}|_{\Gamma_{1}}$.
 Proof.
@@
 Proof.
-This is a special case of Lemma 7.2 applied with $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$ and $\overline{\rho}:\Gamma_{F}\rightarrow G(k),\gamma\mapsto(1,\pi(\gamma))$ via Proposition 5.2 with $\rho_{0}:\Gamma_{F}\rightarrow G(\mathcal{O}),\gamma\mapsto(1,\pi(\gamma))$.
+Let $G=\mathfrak{D}(M)\rtimes\underline{\Delta}$, $\overline{\rho}:\widehat{\Gamma}_{1}\rightarrow G(k),\gamma\mapsto(1,\pi(\gamma))$ and let $\rho_{0}:\widehat{\Gamma}_{1}\rightarrow G(\mathcal{O}),\gamma\mapsto(1,\pi(\gamma))$.
+The map $\Phi\mapsto\Phi\rho_{0}$ induces a natural bijection between $\widehat{Z}^{1}(A)$ and $D^{\square}_{\overline{\rho}}(A)$ for all $A\in\mathfrak{A}_{\mathcal{O}}$; this can be shown by the same argument as in the proof of 5.2.
+Thus $\widehat{Z}^{1}$ is pro-represented by $R^{\square}_{\overline{\rho}}$ and the assertion follows from 5.2, 7.1 and 7.2.
 ∎
@@
 Thus the map $\operatorname{Spec}R^{\square}_{\overline{\rho}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces an $\mathrm{X}(\mu)$-equivariant bijection between the irreducible components.
-Since $A^{\mathrm{gen}}_{G,\overline{\Theta}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$ by 7.13 the map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces a an $\mathrm{X}(\mu)$-equivariant bijection between the sets of irreducible components.
+Since $A^{\mathrm{gen}}_{G,\overline{\Theta}}\cong R^{\mathrm{ps}}_{\overline{\Theta}}[t_{1}^{\pm 1},\ldots,t_{s}^{\pm 1}]$ by 7.13 the map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is $\mathrm{X}(\mu)$-equivariant and induces an $\mathrm{X}(\mu)$-equivariant bijection between the sets of irreducible components.
 The isomorphism $R^{\mathrm{ps}}_{\overline{\Theta}}\cong\mathcal{O}\llbracket H_{1}(\Gamma,M)^{\wedge,p}\rrbracket$ allows us to consider $R^{\mathrm{ps}}_{\overline{\Theta}}$ as an $\mathcal{O}[\mu]$-algebra.
@@
 Proof.
-If $N$ is finitely generated over $\mathbb{Z}$ then the statement is proved in [4, Corollary 2, p.
+If $N$ is finitely generated over $\mathbb{Z}$ then the statement is proved in [5, Corollary 2, p.
 105].
@@
 We then have an isomorphism $\mathfrak{p}_{E}^{n}\otimes M\cong V\otimes M$ of $\Delta$-modules.
-Since $\mathfrak{p}_{E}^{n}\cong\mathcal{O}_{E}$ is isomorphic to $\mathcal{O}_{F}[\Delta]$ as $\mathbb{Z}_{p}[\Delta]$-module, see the proof of [13, Section 1.4], $\mathcal{O}_{E}\otimes M$ is free and thus $(\mathcal{O}_{E}\otimes M)_{\Delta}\cong\mathcal{O}_{F}\otimes M$.
+Since $\mathfrak{p}_{E}^{n}\cong\mathcal{O}_{E}$ is isomorphic to $\mathcal{O}_{F}[\Delta]$ as $\mathbb{Z}_{p}[\Delta]$-module, see the proof of [14, Section 1.4], $\mathcal{O}_{E}\otimes M$ is free and thus $(\mathcal{O}_{E}\otimes M)_{\Delta}\cong\mathcal{O}_{F}\otimes M$.
 Thus $\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}_{p}}((\mathcal{O}_{E}^{\times})^{\wedge,p}\otimes M)_{\Delta}=[F:\mathbb{Q}_{p}]\cdot\mathop{\mathrm{rank}}\nolimits_{\mathbb{Z}}M$ and the assertion follows.
@@
 Proof.
-We have $\mathcal{A}^{\wedge,p}\cong\mathcal{A}\otimes\mathbb{Z}_{p}$ and $(\mathcal{A}^{\Delta})^{\wedge,p}\cong\mathcal{A}^{\Delta}\otimes\mathbb{Z}_{p}$ by [14, Tag 00MA].
+We have $\mathcal{A}^{\wedge,p}\cong\mathcal{A}\otimes\mathbb{Z}_{p}$ and $(\mathcal{A}^{\Delta})^{\wedge,p}\cong\mathcal{A}^{\Delta}\otimes\mathbb{Z}_{p}$ by [15, Tag 00MA].
 We may express $\mathcal{A}^{\Delta}$ as the kernel of
@@
 $0\rightarrow(E^{\times}\otimes M)^{\wedge,p}\rightarrow(\mathcal{E}\otimes M)^{\wedge,p}\rightarrow(I_{\Delta}\otimes M)^{\wedge,p}\rightarrow 0$
-Since $M$ is a free $\mathbb{Z}$-module of finite rank and $(E^{\times})^{\wedge,p}\cong\Gamma^{\mathrm{ab},p}_{E}$ we $(E^{\times}\otimes M)^{\wedge,p}\cong(E^{\times})^{\wedge,p}\otimes M\cong\Gamma^{\mathrm{ab},p}_{E}\otimes M$.
+Since $M$ is a free $\mathbb{Z}$-module of finite rank and $(E^{\times})^{\wedge,p}\cong\Gamma^{\mathrm{ab},p}_{E}$ we have
+$(E^{\times}\otimes M)^{\wedge,p}\cong(E^{\times})^{\wedge,p}\otimes M\cong\Gamma^{\mathrm{ab},p}_{E}\otimes M.$
 Since $I_{\Delta}$ is a free $\mathbb{Z}$-module $(I_{\Delta}\otimes M)^{\wedge,p}\cong I_{\Delta}\otimes M^{\wedge,p}$, so that we obtain an exact sequence of $\mathbb{Z}_{p}[\Delta]$-modules:
@@
 Let $\Gamma_{E}$ be the kernel of the composition $\Gamma_{F}\overset{\overline{\rho}}{\longrightarrow}G(k)\rightarrow(G/G^{0})(k)$ and let $\Delta:=\operatorname{Gal}(E/F)$.
-Let $\pi:\Gamma_{F}\rightarrow\Delta$ be the projection map.
+Let $\pi:\Gamma_{F}\rightarrow\Delta$ and $\Pi:G\rightarrow G/G^{0}$ be the projection maps.
 We let $D^{\mathrm{ps}}_{\overline{\Theta}}:\mathfrak{A}_{\mathcal{O}}\rightarrow\operatorname{Set}$ be the functor
@@
 $0\rightarrow E^{\times}\rightarrow W_{E/F}\rightarrow\Delta\rightarrow 0,$
-corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$ by [15, (1.2)].
+corresponding to the fundamental class $[u_{E/F}]\in H^{2}(\Delta,E^{\times})$ by [16, (1.2)].
 Let $\widehat{W}_{E/F}$ be the profinite completion of $W_{E/F}$.
@@
 Lemma 9.1.
-Every continuous representation $\rho:\Gamma_{F}\rightarrow G(A)$, where $A$ is a finite discrete $\mathcal{O}$-algebra, factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$.
+Every continuous representation $\rho:\Gamma_{F}\rightarrow G(A)$ satisfying $\Pi\circ\rho=\pi$, where $A$ is a finite discrete $\mathcal{O}$-algebra, factors through the quotient $\Gamma_{F}\twoheadrightarrow\widehat{W}_{E/F}$.
 Proof.
@@
 If $L^{\prime}$ is a finite extension of $L$ with the ring of integers $\mathcal{O}^{\prime}$ and residue field $k^{\prime}$ then the functor $D^{\mathrm{ps}}_{\overline{\Theta}_{k^{\prime}}}:\mathfrak{A}_{\mathcal{O}^{\prime}}\rightarrow\operatorname{Set}$ is pro-representable by $R^{\mathrm{ps}}_{\overline{\Theta}}\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$ and analogous statements hold for $X^{\mathrm{gen}}_{G,\overline{\Theta}_{k^{\prime}}}$ and $D^{\square}_{\overline{\rho}_{k^{\prime}}}$.
-Thus it is enough to prove the statement for $\mathcal{O}^{\prime}=\mathcal{O}$, while allowing to replace $L$ by a finite extension.
+Thus it is enough to prove the statement for $\mathcal{O}^{\prime}=\mathcal{O}$, after replacing $L$ by a finite extension.
 Since $G$ is a generalised torus, after replacing $L$ by a finite unramified extension we may assume that $G^{0}$ is split and $G/G^{0}$ is a constant group scheme.
@@
 Lemma 7.10 and (65) imply that $R^{\mathrm{ps}}_{\overline{\Theta}}$ is isomorphic to $\mathcal{O}\llbracket N\rrbracket$, where $N$ is the pro-$p$ completion of $(E^{\times}\otimes M)^{\Delta}$.
-It follows from Lemma 8.7 that $N\cong(\Gamma^{\mathrm{ab}}_{E}\otimes M)^{\Delta}$, which is a finitely generated $\mathbb{Z}_{p}$-module of rank $r$ and torsion subgroup isomorphic to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$ by 8.8.
+It follows from Lemma 8.7 that $N\cong(\Gamma^{\mathrm{ab},p}_{E}\otimes M)^{\Delta}$, which is a finitely generated $\mathbb{Z}_{p}$-module of rank $r$ and torsion subgroup isomorphic to $(\mu_{p^{\infty}}(E)\otimes M)^{\Delta}$ by 8.8.
 This yields the isomorphisms in part (3).
 Part (4) follows from Proposition 7.13.
-Part (5) follows from part (4) and Lemma 7.4 as $E^{\times}/(E^{\times})^{p}$ is finite and the map $\rho\mapsto\Theta_{\rho}$ induces a map of local $\mathcal{O}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$.
+The map $\rho\mapsto\Theta_{\rho}$ induces a map of local $\mathcal{O}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ and hence a map of $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebras $A^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$.
+Since $E^{\times}/(E^{\times})^{p}$ is finite, Lemma 7.4 implies that $R^{\square}_{\overline{\rho}}$ is the completion of $\mathcal{O}(\operatorname{Rep}^{W_{E/F}}_{G,\pi})$ with respect to the maximal ideal corresponding to $\overline{\rho}$.
+By considering the composition $\mathcal{O}(\operatorname{Rep}^{W_{E/F}}_{G,\pi})\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ we deduce that that it induces a natural isomorphism between $R^{\square}_{\overline{\rho}}$ and the completion of $A^{\mathrm{gen}}_{G,\overline{\Theta}}$ with respect to the maximal ideal corresponding to $\overline{\rho}$.
+Part (5) then follows from parts (3) and (4).
 ∎
@@
 Proof.
-Since the map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\mathrm{ps}}_{\overline{\Theta}}$ is faithfully flat, the assertion follows from [14, Tag 06CM] and part (5) of Theorem 9.3.
+Since the map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow\mathcal{O}^{\prime}\otimes_{\mathcal{O}}R^{\mathrm{ps}}_{\overline{\Theta}}$ is faithfully flat, the assertion follows from [15, Tag 06CM] and part (5) of Theorem 9.3.
 ∎
@@
 The map $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A^{\mathrm{gen}}_{G,\overline{\Theta}}$ is formally smooth by the same argument as in Corollary 9.4 using part (4) of Theorem 9.3.
-Since it is of finite type it is smooth by [14, Tag 00TN].
-The map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}:=\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is flat and of finite presentation and hence open by [14, Tag 00I1].
-By [14, Tag 004Z] it is enough to show the fibres of $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ are irreducible.
+Since it is of finite type it is smooth by [15, Tag 00TN].
+The map $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}:=\operatorname{Spec}R^{\mathrm{ps}}_{\overline{\Theta}}$ is flat and of finite presentation and hence open by [15, Tag 00I1].
+By [15, Tag 004Z] it is enough to show the fibres of $X^{\mathrm{gen}}_{G,\overline{\Theta}}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ are irreducible.
 To ease the notation we let $R=R^{\mathrm{ps}}_{\overline{\Theta}}$, $A=A^{\mathrm{gen}}_{G,\overline{\Theta}}$, $R^{\prime}=R\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$ and $A^{\prime}=A\otimes_{\mathcal{O}}\mathcal{O}^{\prime}$.
@@
 Remark 9.7.
-Lemma 9.6 implies that the scheme $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ coincides with the scheme $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ defined in [11] for a generalised reductive group $G$, see [11, Proposition 7.3].
+Lemma 9.6 implies that the scheme $X^{\mathrm{gen}}_{G,\overline{\Theta}}$ coincides with the scheme $X^{\mathrm{gen}}_{G,\bar{\rho}^{\mathrm{ss}}}$ defined in [12] for a generalised reductive group $G$, see [12, Proposition 7.3].
 Corollary 9.8.
@@
 If $\ker f$ is torsion then $\mathcal{P}_{1}^{\mathrm{tf}}$ is a submodule of $\mathcal{P}_{2}^{\mathrm{tf}}$ and the fibre $k\otimes_{R_{1}^{\mathrm{red}}}R_{2}^{\mathrm{red}}$ is isomorphic to $k\llbracket\mathcal{P}^{\mathrm{tf}}_{2}/\mathcal{P}^{\mathrm{tf}}_{1}\rrbracket$.
-Since the dimension of the fibre is equal to $d_{2}-d_{1}$, where $d_{i}=\dim_{\mathbb{Q}_{p}}(\mathcal{P}_{i}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$, we deduce from [14, Tag 00R4] that $R_{1}^{\mathrm{red}}\rightarrow R_{2}^{\mathrm{red}}$ is flat.
+Since the dimension of the fibre is equal to $d_{2}-d_{1}$, where $d_{i}=\dim_{\mathbb{Q}_{p}}(\mathcal{P}_{i}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$, we deduce from [15, Tag 00R4] that $R_{1}^{\mathrm{red}}\rightarrow R_{2}^{\mathrm{red}}$ is flat.
 Let $M$ be the character lattice of $G^{0}$ and let $N$ be the character lattice of $H^{0}$.